### v1

# Open-Set Deepfake Detection: A Parameter-Efficient Adaptation Method with Forgery Style Mixture

Abstract
Open-set face forgery detection poses significant security threats and presents substantial challenges for existing detection models. These detectors primarily have two limitations: they cannot generalize across unknown forgery domains and inefficiently adapt to new data. To address these issues, we introduce an approach that is both general and parameter-efficient for face forgery detection. It builds on the assumption that different forgery source domains exhibit distinct style statistics. Previous methods typically require fully fine-tuning pre-trained networks, consuming substantial time and computational resources. In turn, we design a forgery-style mixture formulation that augments the diversity of forgery source domains, enhancing the model’s generalizability across unseen domains. Drawing on recent advancements in vision transformers (ViT) for face forgery detection, we develop a parameter-efficient ViT-based detection model that includes lightweight forgery feature extraction modules and enables the model to extract global and local forgery clues simultaneously. We only optimize the inserted lightweight modules during training, maintaining the original ViT structure with its pre-trained ImageNet weights. This training strategy effectively preserves the informative pre-trained knowledge while flexibly adapting the model to the task of Deepfake detection. Extensive experimental results demonstrate that the designed model achieves state-of-the-art generalizability with significantly reduced trainable parameters, representing an important step toward open-set Deepfake detection in the wild. The code is available at https://github.com/ChenqiKONG/OSDFD.
Index Terms:
Deepfakes, face forgery detection, open-set, style mixture, generalization, robustness, parameter-efficient learning.
I Introduction
Figure 1: False Positive Rates (FPR) and False Negative Rates (FNR) of Xception, EfficientNet-B4, and ViT-B on four unforeseen Deepfake datasets: DFDC, DFR, WDF, and FFIW.
The rapid advancement of Artificial Intelligence Generated Content (AIGC) has led to significant progress in face forgery techniques, producing increasingly sophisticated fake faces that are indistinguishable from the human eye. Such advancements have fueled the proliferation of multimedia content on social media, posing high risks of disinformation, misinformation, and identity theft [1]. The recent emergence of diffusion and other advanced large vision-language models has exacerbated these challenges [2]. Consequently, developing effective detection methods to combat digital face forgery attacks has become critically important [3].
Face forgery detection has long posed a challenging yet crucial problem. Early methodologies proposed extracting hand-crafted features such as LBP maps [4], color components [5], and DCT maps [6] as informative indicators for Deepfake detection. Subsequent works focused on capturing anomalies like eye-blinking [7], head pose artifacts [8], and face-warping features [9]. Many methods aimed to improve detection accuracy by developing learning-based models using generic architectures such as Xception Net [10], Efficient Net [11], and Capsule Net [12]. Later approaches, including Face X-ray [13], SBI [14], DCL [15], and RECCE [15] sought to capture common forgery features to enhance model generalizability. With the advent of ViT [16], numerous ViT-based detectors [17, 18, 19, 20] have been developed. These approaches exploit global receptive fields and exhibit superior detection performance.
However, previous face forgery detection methods still face two major challenges in real-world applications: (1) Existing methods struggle to generalize in uncontrolled environments. These models may encounter unforeseen forgery domains in the wild, which can lead to significant performance degradation; (2) In practical scenarios, detection models often require fine-tuning newly acquired data to adapt to new domains. However, adapting pre-trained detection models to new data is computationally costly and impedes deployment in real-world settings, especially on mobile devices with limited computing resources.
To address these two challenges, this paper introduces a generalized and parameter-efficient approach for Open-Set DeepFake Detection (OSDFD). Different forgery-style statistics may exhibit unique properties critical for face forgery detection. Fig. 1 shows that existing forgery detectors have a significantly higher false negative rate (FNR) than the false positive rate (FPR) in unforeseen forgery domains, indicating a tendency to misclassify fake faces as real when encountering new domains in uncontrolled environments. Grounded in this observation, we propose a forgery style mixture module that enhances the diversity of source forgery domains by randomly mixing the feature statistics of different forgery styles, thereby effectively improving the model’s generalizability.
Furthermore, most prior art utilizes generic CNNs and ViTs as backbones and fully fine-tune the networks from ImageNet weights. However, fully fine-tuning a network is costly, hindering its real-world applications. To mitigate this issue, we integrate lightweight Adapter and LoRA layers into the ViT backbone. During training, we freeze plain ViT backbones with ImageNet weights and solely update the lightweight Adapter and LoRA layers. These designed modules capture local and global forgery artifacts using slim components. Additionally, these modules preserve substantial ImageNet knowledge and enable flexible model adaptation for face forgery detection. Fig. 2 illustrates the average cross-dataset AUC performance across six unforeseen Deepfake datasets versus the number of activated parameters. This demonstrates that our method achieves superior generalizability with the fewest activated parameters.
This work extends upon our previous conference work [21]. The main distinctions between this work and [21] are summarized as follows: (1) We introduce the Central Difference Convolution (CDC) adapter to incorporate local forgery priors into the expressive ViT model, thereby enhancing the model’s detection performance, (2) We propose a forgery style mixture module that combines source forgery domains to improve the model’s generalizability across unforeseen forgery domains; (3) We further reduce the dimension of LoRA layers, enabling more parameter-efficient face forgery detection; (4) We conduct extensive additional experiments compared with more SOTA methods, including cross-dataset evaluations, robustness experiments, visualization results, and ablation studies.
In summary, the main contributions of this work are as follows:
  • We design a forgery-style mixture module to enhance the diversity of source domains. The model’s generalizability is effectively improved by randomly mixing forgery styles.
  • We introduce a CDC adapter and a LoRA module into pre-trained plain ViT models. This design successfully captures forgery-related features and enables parameter-efficient face forgery detection.
  • We carry out extensive experiments and demonstrate that our model achieves state-of-the-art generalizability and robustness with only 1.34M activated parameters.
Figure 2: Average cross-dataset AUC score across six unseen datasets vs. the number of activated parameters. Our OSDFD method achieves the best generalizability with the fewest trainable parameters.
We arranged the remaining of this paper into four more sections, Sec. II comprehensively reviews previous literature on face forgery detection and parameter-efficient fine-tuning. Sec. III details the proposed detection framework. Sec. IV presents extensive experimental results under various settings—finally, Sec. V concludes the paper and discusses possible future research directions.
II Related Work
II-A Face forgery detection
Early face forgery detection methods initially focused on extracting hand-crafted features, such as eye-blinking frequency [7], head pose inconsistency [8], face warping artifacts [9], and rPPG features [22]. With the rise of deep learning and artificial intelligence, numerous learning-based methods emerged to improve detection accuracy. Some generic CNNs, such as the Xception network [10, 23, 24, 6], EfficientNet-b4 [11, 25], and Capsule network [12], have been utilized for face forgery detection, showing promising results. However, these methods often need to be more balanced in training data, leading to poor generalization capabilities. In response, some approaches have focused on extracting common forgery features. For example, Face X-ray [13] and SBI [14] aimed to extract common face blending features, enhancing model generalizability. F3Net and Luo et al. [24] mined frequency artifacts, while DCL [15] and CFM [25] employed contrastive learning schemes for more generalized Deepfake detection. With the rapid advancements in vision transformers (ViT), several ViT-based approaches [17, 18, 19, 26, 21, 27, 28] have been developed, showing superior detection performance. ICT [17] introduced the Identity Consistency Transformer to capture identity inconsistency within fake faces, while F2Trans designed the High-Frequency Fine-Grained Transformer, achieving more generalized and robust forgery detection. However, the scalability of ViT-based methods is largely constrained by the quadratic complexity of attention computation, limiting their deployment, particularly on mobile devices with limited computing resources. In response, this paper introduces a parameter-efficient fine-tuning (PEFT) strategy for face forgery detection to address this issue. Furthermore, we combine different forgery styles to augment the forgery source domains, enhancing the model’s generalization capability.
II-B Parameter-efficient fine-tuning (PEFT)
The attention mechanism has demonstrated significant success across various computer vision tasks. However, the self-attention in ViT scales quadratically concerning the input patch number, leading to substantial computational demands during training. Several methods have been proposed to address this issue, including parameter-efficient fine-tuning (PEFT). PEFT selectively updates a small portion of the model parameters, facilitating model adaptation to downstream tasks. Adapters [29, 30] are typically inserted between feedforward and normalization layers of transformers, comprising a down-sample and up-sample layer. For example, Low-Rank Adaptation (LoRA) [31] decomposes a matrix into two trainable low-rank matrices, significantly reducing the number of parameters compared to the original matrix. Visual Prompt Tuning (VPT) [32] introduces a small number of trainable parameters in the input space, treating additional tokens as learnable pixels. Convpass [33] aims to inject local priors into expressive ViT models by integrating convolutional operations, enhancing performance across various computer vision tasks. Neural Prompt Search (NOAH) [34] further advances by incorporating Adapter, LoRA, and VPT into vision transformers and optimizing their design through a neural architecture search algorithm. Despite these advancements, the development of PEFT methods for face forgery detection has been largely understudied, yet it is crucial in practical scenarios. In this paper, we integrate lightweight Adapter and LoRA modules into the plain ViT backbone, optimizing only these modules while keeping the ViT backbone fixed with ImageNet weights during training. The designed modules effectively capture global and local face forgery features more efficiently, without the danger of catastrophic forgetting typical of fully finetuning models [35].
III Proposed Method
III-A Overall framework
Fig. 3 illustrates the overview of the proposed face forgery detection method. The designed framework relies upon ViT, which is pre-trained on ImageNet as the backbone. As Fig. 3 (a) depicts, the features extracted by the transformer blocks are fed forward to the Forgery Style Mixture module. This module exclusively mixes the statistics of forgery styles, augmenting the forgery source domains during training. The processed features are delivered to the MLP head for the final decision-making. Fig. 3 (b) illustrates the details of transformer blocks. We insert the trainable Adapter and LoRA layers into the ViT block. The ViT backbone is fixed with pre-trained ImageNet weights. Fig. 3 (c) and Fig. 3 (d) detail the Adapter and LoRA layers. In Fig. 3 (c), the input hidden state $h^{in}$ is followed by three trainable convolutional layers. Moreover, $h^{in}$ is simultaneously multiplied with the ImageNet pretrained weight. The output hidden state $h^{out}$ is the element summation of the two features. In turn, the LoRA layer dedicated in Fig. 3 (d) introduces two trainable low-rank matrices $W^{down}\in\mathbb{R}^{d\times r}$ and $W^{up}\in\mathbb{R}^{r\times d}$, where $r<<d$. Moreover, $h^{out}$ is obtained by conducting an element-wise summation of the two outputs.
Figure 3: Overview of the designed face forgery detection framework. (a) Overall model structure; (b) Structure of the designed transformer block; (c) Details of the designed adapter layer; (d) Details of the designed LoRA layer.
Figure 4: Illustration of the Central-Difference Convolution (CDC) pipeline.
III-B Forgery-aware PEFT
Our proposed method aims to capture critical forgery clues using lightweight modules to facilitate face forgery detection in practical scenarios. We integrate forgery-aware PEFT modules into plain ViTs. These modules include the LoRA layer, injected into the self-attention blocks, and the adapter layer, integrated into the feed-forward networks (FFN).
As Fig. 3 (c) depicts, the input hidden state $h^{in}$ is simultaneously forwarded to the fixed ImageNet MLP weights and the designed trainable convolutional layers. Subsequently, the outputs of these processes are element-wise summed before being passed to the subsequent vision transformer blocks:
$h^{out}={\rm MLP}(h^{in})+{\rm Adapter}(h^{in}).$ (1)
The Adapter layer is formulated as:
${\rm Adapter}(h^{in})={\rm Conv_{1\times 1}^{up}}({\rm CDC}({\rm Conv_{1\times 1}^{down}}(h^{in}))).$ (2)
We first expand the input hidden state $h^{in}$ into a 2-D dimension. Next, we reduce the dimension of the expanded feature using a $1\times 1$ convolutional layer, effectively decreasing the number of parameters activated in the Adapter layer. We then design the Central Difference Convolution (CDC) operator to expose local forgery artifacts for each input query. The pipeline of the CDC operator is illustrated in Fig. 4. The output feature map $x^{out}$ can be calculated by:
$x^{out}=\sum_{i\in\Omega}w_{i}(x^{in}_{i}-x_{c}),$ (3)
Figure 5: Illustration of forgery style mixture pipeline. The designed feature extractor processes real faces and forgery source domain data, and the extracted features are subsequently delivered to the forgery-style mixture module.
where $x_{c}$ indicates the center pixel of the window $\Omega$, and $x_{5}$ represents $x_{c}$ in Fig. 4. $x_{i}$ indicates the peripheral pixels within $\Omega$. $w_{i}$ denotes the corresponding convolutional weights. Finally, the second $1\times 1$ Conv layer restores the channels of the feature, and the produced feature is flattened to the same shape as $h^{in}$. Our designed adapter layer effectively introduces inductive bias into standard ViTs, guiding the model in learning local artifacts for face forgery detection. Additionally, considering the effectiveness of high-frequency features in the task of face forgery detection [6, 36, 37, 38], our CDC operator enables the model to extract abundant high-frequency information from input faces, thereby enhancing detection performance.
In addition to the Adapter layer, which introduces local dependencies into plain ViT backbones, we further design the Low-Rank Adaptation (LoRA) layer to obtain global receptive fields with minimal parameters. As depicted in Fig. 3 (d), the output hidden state $h^{out}$ is calculated by:
$h^{out}=W_{q/k/v}h^{in}+{\rm LoRA}(h^{in}).$ (4)
The $W_{q/k/v}$ parameter denotes the pre-trained ImageNet weights fixed during the training process. Moreover, the LoRA module consists of two low-rank matrices:
${\rm LoRA}(h^{in})=h^{in}W_{q/k/v}^{down}W_{q/k/v}^{up},$ (5)
where $h^{in}\in\mathbb{R}^{d\times d}$, $W_{q/k/v}^{down}\in\mathbb{R}^{d\times r}$, $W_{q/k/v}^{up}\in\mathbb{R}^{r\times d}$, and $r<<d$. We optimize the slim, decomposed low-rank matrices during training instead of the redundant pre-trained weights. As such, the number of trainable parameters is significantly reduced. In this work, we set the hidden state dimension $d$ of ViT-B to 768, and the rank $r$ is defaulted to 8.
Figure 6: Pipeline of the proposed forgery feature mixture process.
The designed forgery-aware PEFT modules bear the following advantages for face forgery detection: (1) they effectively preserve knowledge learned from ImageNet, while adaptively learning forgery-specific information; and (2) they simultaneously capture expressive global features and extract informative local forgery clues, thereby enhancing the final detection performance. These advantages jointly contribute to a more generalized approach to face forgery detection.
III-C Forgery style mixture
In face forgery detection, the training dataset often encompasses multiple forgery techniques. This training setting accommodates practical scenarios where available source domains are utilized. In this vein, we denote the training dataset as $D^{S}=\{R^{S},F_{1}^{S},F_{2}^{S},...,F_{N}^{S}\}$. $R^{S}$ denotes the source real faces, and $F_{i}^{S}$ is the $i_{th}$ source forgery domain, $i\in\{1,2,...,N\}$. As such, the testing data can be denoted as $D^{T}=\{R^{T},F_{1}^{T},F_{2}^{T},...,F_{M}^{T}\}$. Similarly, $R^{T}$ and $F^{T}$ are target real and forgery domains, while $M$ denotes the number of target forgery domains. Our preliminary experiment, illustrated in Fig. 1, reveals a high false negative rate (FNR) and a low false positive rate (FPR) in cross-domain evaluations, suggesting a negligible domain gap between $R^{S}$ and $R^{T}$ and significant domain gaps between $F^{S}$ and $F^{T}$. As such, we need to augment source forgery domains to mitigate this issue. Fig. 5 depicts the proposed forgery-style mixture pipeline. The training batch $X^{S}=\{R^{S},F_{1}^{S},F_{2}^{S},...,F_{N}^{S}\}$ is processed by our designed feature extractor, which shares weights across all source domains. The extracted features from the source forgery domains are combined in the designed forgery-style mixture module, and the mixed features are forwarded to the classifier to perform forgery detection.
Next, we elaborate on the details of the devised forgery style mixture module, as depicted in Fig. 6. This module processes a training batch of real and forged face images sampled from $N$ source domains, where $N=3$, for instance, is shown in Fig. 6. Green squares denote real faces in this representation, while red circles indicate the forgery faces. The input feature $f^{in}$ proceeds to our designed Forgery Style Mixture module.
We first sort $f^{in}$ to $f^{sort}$ according to the domain labels, where $f_{R}$ represents real features, and $f_{Fi}$ with $i\in\{1,2,3\}$, denotes features from distinct forgery domains. The superscript of each sample indicates the feature index of the domain. To augment the forgery style, we first shuffle the fake features of $f^{sort}$, ensuring that each forgery domain label of the shuffled $\widetilde{f}^{sort}$ differs from the original $f^{sort}$. In addition, we keep the real feature order unchanged. Subsequently, we blend the forgery feature styles of the original $f^{sort}$ and the shuffled $\widetilde{f}^{sort}$.
We draw inspiration from Adaptive Instance Normalization (AdaIN) [39], a widely employed image style transfer technique. The source style of $p$ can be transferred to the target style of $q$ by:
${\rm AdaIN}(p)=\sigma(q)\frac{p-\mu(p)}{\sigma(p)}+\mu(q),$ (6)
where $\mu(p)$ and $\sigma(p)$ are the spatial mean and standard deviation of the feature tensor $p$.
Inspired by the AdaIN, which utilizes the feature statistics of a target style input to transfer the image style, we first calculate the mixed feature statistics of $f$ and $\tilde{f}$:
$\gamma_{mix}=\delta\sigma(f)+(1-\delta)\sigma(\tilde{f}),$ (7)
$\eta_{mix}=\delta\mu(f)+(1-\delta)\mu(\tilde{f}),$ (8)
where $\delta$ is a weight sampled from Beta distribution, i.e., $\delta\sim Beta(0.1,0.1)$. $\gamma_{mix}$ and $\eta_{mix}$ are the weighted summation of the forgery style statistics of $f$ and $\tilde{f}$. Then, the style mixture is performed by the mixed feature statistics $\gamma_{mix}$ and $\eta_{mix}$ to the original feature $f$:
$f_{mix}={\rm MixStyle}(f)=\gamma_{mix}\frac{f-\mu(f)}{\sigma(f)}+\eta_{mix}.$ (9)
Similar to [40], we set a default probability of 0.5 to activate the forgery-style mixture module. Finally, we restore the original index order of the mixed features $f^{sort}$ to obtain $f^{out}$, which is forwarded to the classifier for final decision-making. In the inference stage, the forgery-style mixture is deactivated.
During training, the designed forgery-style mixture module augments the source forgery domains, effectively mitigating the overfitting problem and enhance the model’s generalizability. This design enhances the model’s generalization for open-set face forgery detection.
TABLE I: High quality (c23) cross-manipulation detection performance on unseen forgery methods. The trainable parameter number of each model is reported to evaluate the training efficiency.
[…]
AUC ACC AUC ACC AUC ACC AUC ACC AUC ACC
ResNet18 [41] CVPR16 11.18M 0.813 0.656 0.746 0.596 0.464 0.476 0.688 0.528 0.678 0.564
Xception [10] CVPR17 20.81M 0.907 0.795 0.753 0.558 0.460 0.472 0.744 0.557 0.716 0.596
EN-B4 [11] ICML19 19.34M 0.485 0.495 0.556 0.523 0.517 0.517 0.493 0.500 0.513 0.509
All-train EN-B4 [11] ICML19 19.34M 0.911 0.824 0.801 0.633 0.543 0.500 0.774 0.608 0.757 0.641
Focal-loss EN-B4 [11] ICML19 19.34M 0.903 0.813 0.798 0.608 0.503 0.484 0.759 0.604 0.741 0.627
Forensics Transfer [42] ArXiv18 - - 0.720 - 0.645 - 0.460 - 0.569 - 0.599
Multi-task [43] BTAS19 - - 0.703 - 0.587 - 0.497 - 0.603 - 0.598
MLDG [44] AAAI18 62.38M 0.918 0.842 0.771 0.634 0.609 0.527 0.780 0.621 0.770 0.656
LTW [45] AAAI21 20.37M 0.927 0.856 0.802 0.656 0.640 0.549 0.773 0.653 0.786 0.679
ViT-B [16] ICLR21 85.80M 0.771 0.701 0.656 0.582 0.510 0.498 0.554 0.517 0.623 0.575
CFM [46] TIFS24 25.37M 0.880 0.617 0.814 0.557 0.631 0.516 0.644 0.508 0.742 0.550
SCLoRA [21] MIPR23 4.72M 0.935 0.862 0.875 0.753 0.651 0.554 0.707 0.626 0.792 0.699
OSDFD Ours 1.34M 0.962 0.885 0.932 0.846 0.698 0.595 0.781 0.682 0.843 0.752
TABLE II: Low quality (c40) cross-manipulation detection performance on unseen forgery methods. The trainable parameter number of each model is reported to evaluate the training efficiency.
[…]
AUC ACC AUC ACC AUC ACC AUC ACC AUC ACC
ResNet18 [41] CVPR16 11.18M 0.741 0.673 0.648 0.600 0.634 0.594 0.598 0.567 0.655 0.609
Xception [10] CVPR17 20.81M 0.766 0.694 0.696 0.643 0.626 0.593 0.597 0.552 0.671 0.621
EN-B4 [11] ICML19 19.34M 0.451 0.485 0.537 0.505 0.512 0.503 0.499 0.497 0.500 0.498
All-train EN-B4 [11] ICML19 19.34M 0.753 0.676 0.674 0.614 0.614 0.580 0.600 0.564 0.660 0.609
Focal-loss EN-B4 [11] ICML19 19.34M 0.749 0.674 0.672 0.610 0.596 0.575 0.605 0.566 0.656 0.606
Forensics Transfer [42] ArXiv18 - - 0.682 - 0.550 - 0.530 - 0.550 - 0.578
Multi-task [43] BTAS19 - - 0.667 - 0.565 - 0.517 - 0.560 - 0.577
MLDG [44] AAAI18 62.38M 0.730 0.671 0.617 0.581 0.617 0.581 0.607 0.569 0.643 0.601
LTW [45] AAAI21 20.37M 0.756 0.691 0.724 0.657 0.681 0.625 0.608 0.585 0.692 0.640
ViT-B [16] ICLR21 85.80M 0.739 0.643 0.650 0.595 0.592 0.560 0.552 0.538 0.633 0.584
CFM [46] TIFS24 25.37M 0.789 0.686 0.622 0.541 0.688 0.607 0.566 0.517 0.666 0.588
SCLoRA [21] MIPR23 4.72M 0.818 0.735 0.686 0.638 0.710 0.653 0.582 0.542 0.699 0.642
OSDFD Ours 1.34M 0.853 0.762 0.717 0.663 0.713 0.661 0.630 0.595 0.728 0.670
III-D Objective function
During training, as we augment source-forgery features using the designed forgery style mixture module, we must train a classifier that can effectively distinguish between real face features and the augmented forgery features. Therefore, this work adopts the Single-Center Loss (SCL) [45] to obtain a clearer boundary between bonafide and fake faces. The overall objective function consists of two losses:
$L=L_{BCE}+\lambda L_{SCL},$ (10)
where $\lambda$ balances the two loss components. $L_{BCE}$ represents the binary cross-entropy loss. The Single-Center Loss, denoted as $L_{SCL}$, aims to compact the feature distribution of real faces while simultaneously moving fake features away from the center of real features:
$L_{SCL}=Dist_{R}+max(Dist_{R}-Dist_{F}+margin,0)$ (11)
where $Dist_{R}$ and $Dist_{F}$ denote the average distance between the real feature center $C$ and real feature $f_{R}$ and fake feature $f_{F}$, respectively:
$Dist_{R}=\frac{1}{N_{R}}\sum_{{i=1}}^{N_{R}}||f_{R}^{i}-C||_{2};\ Dist_{F}=\frac{1}{N_{F}}\sum_{{j=1}}^{N_{F}}||f_{F}^{j}-C||_{2}$ (12)
where the real center $C$ is the average of all $N_{R}$ real features $f_{R}$:
$C=\frac{1}{N_{R}}\sum_{{i=1}}^{N_{R}}f_{R}^{i}.$ (13)
We select the features after the second-to-last fully-connected layer to calculate $L_{SCL}$. This loss enhances the discriminative and separable characteristics of the real and fake face features, thereby improving the overall performance of face forgery detection.
IV Experiments and Results
IV-A Experimental setup
IV-A1 Implementation details
In this study, we employ the popular dlib face detector [47] to crop face regions. The proposed framework is implemented on the PyTorch [48] platform, and the model is trained using Adam optimizer [49] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999. We set the learning rate at 3e-5, and no learning rate decay is applied throughout the whole training process. Our default architecture employs the ViT-B as the backbone, and we train the model on a single 2080Ti GPU with a batch size of 48. The training comprises 30,000 iterations, and the ViT-B parameters are initialized with ImageNet-21k [50] weights.
IV-A2 Datasets
We adopt the FaceForensics++ (FF++) dataset [51] for training, which includes original real videos (OR) and four types of forgeries: Deepfakes (DF) [52], Face2Face (FF) [53], FaceSwap (FS) [54], and NeuralTextures (NT) [55]. Each category comprises 1,000 videos in raw, high-quality (c23), and low-quality (c40) formats. Following the official data split strategy [51], we allocate 720 videos for training, 140 for validation, and 140 for testing. Raw videos represent non-compressed videos, whereas c23 and c40 videos denote slightly compressed and highly compressed videos. Our experiments exclusively employ the compressed c23 and c40 videos to align with practical applications.
IV-A3 Evaluation Metrics
To facilitate head-to-head comparisons with previous methodologies, we employ Accuracy (ACC), Area Under the receiver operating characteristic Curve (AUC), and Equal Error Rate (EER) as the evaluation metrics in this work. ACC is a straightforward metric that quantifies the proportion of correctly classified samples from the total number of samples. AUC measures the Area under the Receiver Operating Characteristic (ROC) curve. EER denotes the point of ROC where False Positive Rate (FPR) is equal to the True Positive Rate (TPR). Unless specified otherwise, we report the image-level detection results throughout this paper.
IV-B Cross-manipulation evaluation
In real-world applications, face forgery detectors inevitably encounter data generated by new forgery techniques, leading to significant performance drops. To evaluate the generalization capability, we train the models on the training sets comprising three source manipulation types of FF++. Subsequently, we evaluate the trained models on the test sets of the unseen target manipulation type. Table I presents the detection AUC and ACC of different models on the FF++ high-quality (c23) dataset. We bold the best detection results and underline the second-best results. It can be observed that OSDFD achieves SOTA AUC and ACC performances across all experiments.
Moreover, generalization to poor-quality data is crucial for real-world face forgery detectors. Therefore, we conduct the cross-manipulation experiments on low-quality (c40) data, as shown in Table II, where OSDFD outperforms previous approaches by a significant margin. This underscores the proposed method’s outstanding generalizability across diverse data qualities. Additionally, our approach demonstrates substantial improvements over the ViT-B baseline, demonstrating the effectiveness of the designed framework. Furthermore, compared to our conference version SCLoRA [21], our method consistently achieves superior performance on both c23 and c40 data, which can be attributed to the effectiveness of the designed forgery style mixture module and local CDC feature extractor. Finally, OSDFD comprises the fewest 1.34M trainable parameters, facilitating its deployment in real-world applications, particularly on mobile devices.
TABLE III: Cross-dataset evaluation on six unseen datasets. ‘*’ indicates the trained model provided by the authors. ‘†’ indicates our re-implementation using the public official code. Methods highlighted in blue denote video-level results. The trainable parameter number of each model is reported to evaluate the training efficiency.
Method Venue #Params CDF WDF DFDC-P DFDC DFR FFIW Avgerage
AUC EER AUC EER AUC EER AUC EER AUC EER AUC EER AUC EER
Face X-ray [13] CVPR20 41.97M 74.20 - - - 70.00 - - - - - - - - -
GFF [24] CVPR21 53.25M 75.31 32.48 66.51 41.52 71.58 34.77 - - - - - - - -
LTW [45] AAAI21 20.37M 77.14 29.34 67.12 39.22 74.58 33.81 - - - - - - - -
F2Trans-S [28] TIFS23 117.52M 80.72 - - - 71.71 - - - - - - - - -
SBI* [14] CVPR22 19.34M 81.33 26.94 67.22 38.85 79.87 28.26 68.45 37.23 84.90 23.13 81.47 27.33 77.21 30.29
DCL* [15] AAAI22 19.35M 81.05 26.76 72.95 35.73 71.49 35.90 72.69 34.91 92.26 14.81 82.08 27.75 78.75 29.31
Xception†[10] ICCV19 20.81M 64.14 39.77 68.90 38.67 69.56 36.94 64.52 39.29 91.93 15.52 77.58 31.21 72.77 33.57
RECCE†[56] CVPR22 25.83M 61.42 41.71 74.38 32.64 64.08 40.04 67.61 36.98 92.93 14.74 85.84 24.12 74.38 31.71
EN-B4†[11] ICML19 19.34M 65.24 39.41 67.89 37.21 67.96 37.60 63.48 40.14 92.18 15.51 70.03 35.40 71.13 35.40
CFM [46] TIFS24 25.37M 82.78 24.74 78.39 30.79 75.82 31.67 68.28 37.14 95.18 11.87 78.71 29.28 79.86 27.58
ViT-B [16] ICLR21 85.80M 72.35 34.50 75.29 33.40 75.58 32.11 71.48 35.54 80.47 26.97 68.49 37.12 73.94 33.27
SCLoRA [21] MIPR23 4.72M 73.56 32.87 77.91 29.56 77.08 30.37 72.32 34.52 94.78 11.76 81.42 26.59 79.51 26.59
OSDFD Ours 1.34M 83.35 24.47 78.53 28.46 76.28 31.17 75.52 31.99 95.54 9.90 84.61 24.09 82.31 25.01
Lisiam [57] TIFS22 - 78.21 - - - - - - - - - - - - -
F3Net [6] ECCV20 42.53M 68.69 - - - 67.45 - - - - - - - - -
FTCN [58] ICCV21 26.60M 86.90 - - - 74.00 - - - - - - - - -
SBI* [14] CVPR22 19.34M 88.61 19.41 70.27 37.63 84.80 25.00 71.70 35.27 90.04 17.53 86.34 23.45 81.96 26.38
DCL* [15] AAAI22 19.35M 88.24 19.12 76.87 31.44 77.57 29.55 75.03 30.94 97.41 9.96 86.26 24.32 83.56 24.22
RECCE†[56] CVPR22 25.83M 69.25 34.38 76.99 30.49 66.90 39.39 70.26 34.25 97.15 9.29 92.05 17.74 78.77 27.59
CFM [46] TIFS24 25.37M 89.65 17.65 82.27 26.80 80.22 27.48 70.59 35.02 97.59 9.04 83.81 24.49 84.02 23.41
ViT-B [16] ICLR21 85.80M 78.12 30.59 78.95 29.59 79.43 28.62 73.35 33.05 86.42 19.90 70.34 35.45 77.77 29.53
SCLoRA [21] MIPR23 4.72M 81.11 26.32 82.78 24.85 80.20 26.68 74.96 35.01 98.53 6.63 85.59 22.70 83.86 23.70
OSDFD Ours 1.34M 90.81 16.94 81.42 24.90 79.49 28.31 78.74 29.37 98.89 4.48 90.01 17.50 86.56 20.25
IV-C Cross-dataset evaluation
Due to the significant domain gaps between training and testing datasets [59, 60] in face forgery detection, cross-dataset evaluation has been a persistent challenge. We further conduct cross-destatet evaluations to assess the generalizability of forgery detectors. We train the models on the FF++ (c23) training set (OR, DF, FF, FS, and NT) and test them on six popular Deepfake datasets, including CelebDF-v2 (CDF) [61], WildDeepfake (WDF) [62], DeepFake Detection Challenge (DFDC) [63], DeepFake Detection Challenge Preview (DFDC-P) [63], DeepForensics-1.0 (DFR) [64], and Face Forensics in the Wild (FFIW) [65]. These testing datasets simulate real-world scenarios well, encompassing various environmental variables. AUC and EER performance are reported in Table III, where ‘*’ indicates the trained model provided by the authors, and ‘†’ indicates our re-implementation using the public official code. The upper table presents the frame-level detection results, while the blue-highlighted bottom table shows video-level results. Video-level detection scores are calculated by averaging all frame scores within each video. The best and second-best results are highlighted in bold and underlined, respectively. The trainable parameter number of each detector is provided in the table for reference. Our method, OSDFD, achieves the highest average detection performance in both frame-level and video-level evaluation, with only 1.34M activated parameters. Compared to the ViT-B backbone, OSDFD consistently demonstrates significant performance improvements across all six datasets, demonstrating its exceptional generalization capability. Furthermore, OSDFD outperforms our conference version SCLoRA by a significant margin, underscoring the effectiveness of the designed CDC Adapter layer and forgery-style mixture module.
IV-D Robustness evaluation
Videos circulating on social media platforms often undergo various distortions during transmission. To simulate real-world scenarios, we introduce six common perturbations: Gaussian blur, pink noise, white noise, block, brightness, and darkness. Fig. 8 showcases both the original and corresponding distorted face images. We incorporate five severity levels for each perturbation type to mimic unpredictable environments. The robustness evaluation results are depicted in Fig. 7. OSDFD does not employ data augmentations during training, facilitating efficient training, or fine-tuning for real-world applications. All models depicted in Fig. 7 are trained on the FF++ c23 training set, with distortions applied to the corresponding test set. As the severity of distortion increases, all face forgery detection methods experience performance degradation. Compared to previous arts, our proposed approach achieves superior or comparable AUC performance across all six types of perturbation. Furthermore, the average robustness results across all distortions highlight that OSDFD, with the fewest trainable parameters, demonstrates the best robustness against practical image/video perturbations.
Figure 7: Robustness to common perturbations at five severity levels: Gaussian Blur, Pink Noise, White Noise, Block, Brightness, and Darkness.
Figure 8: Input face image and its perturbated versions.
TABLE IV: Ablation Experiments on different model components. Average Cross-Manipulation detection performance on FF++ c23 dataset.
ViT LoRA Adapter MixStyle SCL AUC ACC
- - - - 0.623 0.575
- - - 0.816 0.719
- - - 0.825 0.723
- - 0.832 0.740
- 0.836 0.743
0.843 0.752
IV-E Ablation study
In this subsection, we carry out extensive ablation experiments to showcase the effectiveness of the designed components, the flexibility of the proposed learning scheme, the rationale behind the adopted training strategy, and the impacts of the mixed features.
IV-E1 Effectiveness of the designed modules
We conduct ablation experiments to study the impact of each designed module. Table IV reports the average cross-manipulation (c23) detection AUC and ACC scores across four experimental trials. The designed LoRA and Adapter modules significantly enhance the model’s generalizability. The combination of LoRA and Adapter enables the model to simultaneously explore global and local anomalies within input forgery face images, enhancing the final detection performance. The forgery-style mixture module further boosts the AUC and ACC scores by augmenting the forgery source domains during training. The Single-Center Loss (SCL) encourages the real features to cluster more tightly while pushing features from fake images away from the real center. This design leads to a clearer decision boundary between real and fake faces, further enhancing the generalizability of the detection. Overall, the experimental results presented in Table IV underscore the effectiveness of the designed components.
The proposed method showcases its adaptability by being flexibly adapted to other ViT backbones, a practical feature for researchers, engineers, and practitioners in the field of computer vision and image processing. Herein, we assess the face forgery detection performance of our OSDFD on various ViT backbones, including ViT-Tiny (ViT-T), ViT-Small (ViT-S), and ViT-Large (ViT-L). The cross-manipulation (c23) AUC results are presented in Table V. It can be readily observed that the average cross-manipulation performance improves with increasing model capacity. Nevertheless, larger ViT models tend to require significantly more computing resources and inference time, thereby impeding their real-world deployment.
In contrast, OSDFD consistently enhances the generalizability of ViT backbones across different model complexities, underscoring its practicality and adaptability. Specifically, our method achieves average AUC enhancements of 11.7%, 15.7%, and 16.7%, respectively. These results demonstrate the capability of OSDFD to seamlessly integrate into existing ViT models in a plug-and-play manner.
In this work, we adopt a Parameter-Efficient Fine-Tuning (PEFT) strategy, a reliable approach, to train the face forgery detection model. Specifically, we only optimize the Adapter and LoRA layers while keeping the other ViT parameters fixed with ImageNet weights. We compare the PEFT strategy with the full fine-tuning (FT) strategy to validate its effectiveness. Table VI presents the cross-manipulation (c23) AUC results, where DF, FF, FS, and NT denote the testing manipulation types. The models are trained on three manipulations and tested on the remaining one. The two experimental settings employ two training strategies while maintaining the same network architectures. The results in Table VI demonstrate that the PEFT approach outperforms the fully FT strategy across all experimental trials. Notably, the average AUC score of PEFT exhibits a substantial enhancement, increasing from 68.2% to 84.3%. Hence, it can be concluded that the improvements in face forgery detection performance benefit from the PEFT strategy adopted. Furthermore, the PEFT strategy effectively preserves the abundant ImageNet prior knowledge while learning the forgery-aware knowledge through the designed Adapter and LoRA modules in an adaptive manner, reinforcing the reliability of our research.
TABLE V: Cross-Manipulation AUC performance on FF++ (c23) dataset with different ViT backbones.
Backbone OSDFD DF FF FS NT AVG
ViT-T - 0.853 0.658 0.488 0.584 0.646
ViT-T 0.919 0.860 0.563 0.709 0.763
ViT-S - 0.870 0.688 0.465 0.566 0.647
ViT-S 0.929 0.897 0.600 0.791 0.804
ViT-L - 0.850 0.708 0.572 0.593 0.681
ViT-L 0.955 0.941 0.710 0.785 0.848
TABLE VI: Cross-Manipulation AUC performance on FF++ (c23) dataset with different training strategies.
Strategy DF FF FS NT AVG
Fully FT 0.896 0.731 0.496 0.603 0.682
PEFT (Ours) 0.962 0.932 0.698 0.781 0.843
Figure 9: Average cross-manipulation ACC scores across four experimental trials with respect to different mixed feature ratios.
Figure 10: Grad-CAM [66] visualization results of the ViT-B baseline and OSDFD on six unforeseen Deepfake datasets.
IV-E2 Impacts of different ratios of mixed features
Our proposed method utilizes a feature mixture module to augment forgery domains in the feature space. As illustrated in Fig. 9, the $x$-axis represents the ratio of the mixed forgery features among all forgery features. We analyze the impact of this ratio by plotting the cross-manipulation (c23) ACC scores against ratio values in Fig. 9. Similarly, we present the average score across four cross-manipulation evaluation trials. A ratio of ‘0.0’ indicates the pure use of mixed forgery features during training, while ‘1.0’ means no forgery-style mixture is applied. Notably, relying solely on original or mixed forgery features results in limited face forgery detection accuracy. In contrast, combining original and mixed forgery features enhances the model’s generalization capability by augmenting the source forgery domains. The model achieves the highest average ACC score when the ratio of mixed forgery features is 50%. This observation underscores the effectiveness of the designed forgery feature mixture module from another perspective.
IV-F Visualization results
In this subsection, we provide visualization results to illustrate further the effectiveness of the proposed face forgery detection method.
IV-F1 t-SNE feature map visualization
In Fig. 11, we present the t-SNE [67] feature embedding visualization results of both the ViT-B baseline and our method OSDFD. Fig. 11 (a) and (b) showcase the results on the unseen CDF dataset, while Fig. 11 (c)-(h) depict the results on the WDF, DFDC, and FFIW datasets, respectively. Red and blue dots represent real and fake data samples in these visualizations. The manifolds of real and fake feature distributions entangle for the baseline model. This observation indicates that the baseline model struggles to discriminate forgery faces from the real ones. In contrast, OSDFD effectively separates real and fake images in the feature space. This improvement underscores the enhanced generalizability achieved by our model, attributed to the designed forgery-style mixture module and forgery-aware PEFT modules.
Figure 11: t-SNE feature embedding visualization of (a) the ViT-B baseline on the CDF dataset, (b) OSDFD on the CDF dataset, (c) the ViT-B baseline on the WDF dataset, (d) OSDFD on the WDF dataset, (e) the ViT-B baseline on the DFDC dataset, (f) OSDFD on the DFDC dataset, (g) the ViT-B baseline on the FFIW dataset, and (h) OSDFD on the FFIW dataset. The red and blue dots represent real and fake data samples, respectively.
IV-F2 Grad-CAM map visualization
We visualize the Grad-CAM [66] maps to identify attention regions of the developed face forgery detector for input queries, as shown in Fig. 10. These maps illustrate the results of the ViT-B baseline model and OSDFD across six unforeseen Deepfake datasets: CDF, WDF, DFDC, DFDC-P, DFR, and FFIW. These datasets encompass practical environmental scenarios, including lighting conditions, head poses, image quality, and facial occlusion. Compared to the baseline model, our proposed method accurately highlights informative manipulated regions. The baseline model sometimes focuses on irrelevant non-forgery areas, and these ‘non-interesting’ regions can lead to wrong decisions. Conversely, our method effectively identifies regions with significant forgery artifacts, contributing to superior face forgery detection performance. Specifically, our approach consistently demonstrates outstanding gradient-based localization across diverse environmental variables, such as extreme head poses in WDF, abnormal color hues in DFDC, facial occlusion (glasses) in DFDC-P, and extreme facial area ratios in FFIW. These observations further underscore the outstanding generalizability of our method in real-world applications.
V Conclusions and Future Work
This paper introduces a parameter-efficient and generalized approach towards open-set face forgery detection. Our method, OSDFD, incorporates lightweight LoRA and Adapter layers into ViT backbones, updating only these modules during training to learn global and local forgery-aware features efficiently. On the other hand, we propose a novel forgery-style mixture module that augments the forgery source domains. In this vein, our method achieves SOTA generalizability and robustness in face forgery detection. An ablation study further underscores that the designed network modules and objective functions contribute to the final detection performance. Our proposed method, OSDFD, can be flexibly adapted to other backbones in a plug-and-play manner. Further ablation experiments demonstrate that the adopted parameter-efficient training strategy is effective in face forgery detection. Our proposed method will shed light on developing future face forgery detectors for real-world applications.
While our proposed method is effective, we limit our scope herein to non-sequential Deepfake detection. Adapting our approach to leverage abundant sequential information could significantly improve video-level face forgery detection. The designed components can be seamlessly integrated into video transformers, achieving superior performance in uncontrolled environments. This path can be explored in future work.

### v2

@@
 To address these issues, we introduce an approach that is both general and parameter-efficient for face forgery detection.
-It builds on the assumption that different forgery source domains exhibit distinct style statistics.
-Previous methods typically require fully fine-tuning pre-trained networks, consuming substantial time and computational resources.
-In turn, we design a forgery-style mixture formulation that augments the diversity of forgery source domains, enhancing the model’s generalizability across unseen domains.
+Our method builds on the assumption that different forgery source domains exhibit distinct style statistics.
+Specifically, we design a forgery-style mixture formulation that augments the diversity of forgery source domains, enhancing the model’s generalizability across unseen domains.
+In addition, previous methods typically require fully fine-tuning pretrained networks, consuming substantial time and computational resources.
 Drawing on recent advancements in vision transformers (ViT) for face forgery detection, we develop a parameter-efficient ViT-based detection model that includes lightweight forgery feature extraction modules and enables the model to extract global and local forgery clues simultaneously.
@@
 Extensive experimental results demonstrate that the designed model achieves state-of-the-art generalizability with significantly reduced trainable parameters, representing an important step toward open-set Deepfake detection in the wild.
-The code is available at https://github.com/ChenqiKONG/OSDFD.
 Index Terms:
@@
 The rapid advancement of Artificial Intelligence Generated Content (AIGC) has led to significant progress in face forgery techniques, producing increasingly sophisticated fake faces that are indistinguishable from the human eye.
-Such advancements have fueled the proliferation of multimedia content on social media, posing high risks of disinformation, misinformation, and identity theft [1].
+Such advancements have fueled the proliferation of multimedia content on social media, posing high risks of disinformation, misinformation, and identity theft [1, 2].
 The recent emergence of diffusion and other advanced large vision-language models has exacerbated these challenges [3].
@@
 However, previous face forgery detection methods still face two major challenges in real-world applications: (1) Existing methods struggle to generalize in uncontrolled environments.
-These models may encounter unforeseen forgery domains in the wild, which can lead to significant performance degradation; (2) In practical scenarios, detection models often require fine-tuning newly acquired data to adapt to new domains.
-However, adapting pre-trained detection models to new data is computationally costly and impedes deployment in real-world settings, especially on mobile devices with limited computing resources.
+These models may encounter unforeseen forgery domains in the wild, which can lead to significant performance degradation; (2) In real-world applications, detection models typically need fine-tuning with newly acquired data to adapt to different domains.
+However, this adaptation process for pre-trained detection models can be computationally expensive, making it challenging to deploy them effectively, particularly on mobile devices with limited computational resources.
+Figure 2: Prediction score distributions of real and fake faces of four unseen Deepfake datasets: DFDC, DFR, WDF, and FFIW.
+The three rows represent three Deepfake detectors: Xception, Efficient Net, and ViT-B.
+It can be observed that domain gaps between different Deepfake datasets primarily impact the detection of forgery faces rather than real faces.
 To address these two challenges, this paper introduces a generalized and parameter-efficient approach for Open-Set DeepFake Detection (OSDFD).
@@
 1 shows that existing forgery detectors have a significantly higher false negative rate (FNR) than the false positive rate (FPR) in unforeseen forgery domains, indicating a tendency to misclassify fake faces as real when encountering new domains in uncontrolled environments.
-Grounded in this observation, we propose a forgery style mixture module that enhances the diversity of source forgery domains by randomly mixing the feature statistics of different forgery styles, thereby effectively improving the model’s generalizability.
+In Fig.
+2, we further demonstrate that domain gaps between different Deepfake datasets primarily impact the detection of forgery faces rather than real faces.
+Based on this observation, we propose a forgery style mixture module to increase the diversity of source forgery domains during training, thereby effectively mitigating forgery domain gaps in cross-dataset evaluations.
+Specifically, Fig.
+2 presents cross-dataset experimental results for three representative face forgery detectors: Xception [11], EfficientNet [12], and ViT-B [17].
+These models are trained on FF++ [22] and evaluated on DFDC [23], DFR [24], WDF [25], and FFIW [26].
+In Fig.
+2, green and red indicate real and fake faces, corresponding to ground-truth labels of ‘0’ and ‘1’, respectively.
+The $x$-axis and $y$-axis represent model prediction scores and the number of corresponding faces.
+As shown in Fig.
+2, while these detectors perform well in identifying real faces, they struggle with fake faces in cross-dataset evaluations.
+This finding suggests that discrepancies between source and target datasets primarily arises from forgery domain gaps.
+Motivated by this observation, we design the forgery style mixture module to enhance the diversity of source forgery domains, thereby improving the model’s generalizability.
 Furthermore, most prior art utilizes generic CNNs and ViTs as backbones and fully fine-tune the networks from ImageNet weights.
@@
 Additionally, these modules preserve substantial ImageNet knowledge and enable flexible model adaptation for face forgery detection.
+While parameter-efficient fine-tuning (PEFT) has been widely adopted in various domains, our work makes a notable contribution by integrating PEFT with a forgery-aware perspective, specifically tailored for open-set face forgery detection—a task that presents unique challenges in generalizability and adaptation.
 Fig.
@@
 The main distinctions between this work and [27] are summarized as follows: (1) We introduce the Central Difference Convolution (CDC) adapter to incorporate local forgery priors into the expressive ViT model, thereby enhancing the model’s detection performance, (2) We propose a forgery style mixture module that combines source forgery domains to improve the model’s generalizability across unforeseen forgery domains; (3) We further reduce the dimension of LoRA layers, enabling more parameter-efficient face forgery detection; (4) We conduct extensive additional experiments compared with more SOTA methods, including cross-dataset evaluations, robustness experiments, visualization results, and ablation studies.
+Figure 3: Average cross-dataset AUC score across six unseen datasets vs. the number of activated parameters. Our OSDFD method achieves the best generalizability with the fewest trainable parameters.
 In summary, the main contributions of this work are as follows:
   • We design a forgery-style mixture module to enhance the diversity of source domains. The model’s generalizability is effectively improved by randomly mixing forgery styles.
-  • We introduce a CDC adapter and a LoRA module into pre-trained plain ViT models. This design successfully captures forgery-related features and enables parameter-efficient face forgery detection.
-  • We carry out extensive experiments and demonstrate that our model achieves state-of-the-art generalizability and robustness with only 1.34M activated parameters.
-Figure 2: Average cross-dataset AUC score across six unseen datasets vs. the number of activated parameters. Our OSDFD method achieves the best generalizability with the fewest trainable parameters.
+  • Our approach explicitly captures forgery-related artifacts by integrating a CDC adapter and a LoRA module into a ViT-based detection framework.
+This targeted design enhances the model’s ability to extract both global and local forgery cues, which is crucial for Deepfake detection.
+  • Our designed components jointly contribute to an effective and scalable solution for open-set Deepfake detection.
+Extensive experiments demonstrate that our model achieves state-of-the-art generalizability and robustness with only 1.34M activated parameters.
 We arranged the remaining of this paper into four more sections, Sec.
@@
 With the rise of deep learning and artificial intelligence, numerous learning-based methods emerged to improve detection accuracy.
-Some generic CNNs, such as the Xception network [10, 23, 24, 6], EfficientNet-b4 [11, 25], and Capsule network [12], have been utilized for face forgery detection, showing promising results.
+Some generic CNNs, such as the Xception network [11, 29, 30, 7, 31], EfficientNet-b4 [12, 32, 33, 34], and Capsule network [13], have been utilized for face forgery detection, showing promising results.
 However, these methods often need to be more balanced in training data, leading to poor generalization capabilities.
@@
 [30] mined frequency artifacts, while DCL [16] and CFM [32] employed contrastive learning schemes for more generalized Deepfake detection.
-With the rapid advancements in vision transformers (ViT), several ViT-based approaches [17, 18, 19, 26, 21, 27, 28] have been developed, showing superior detection performance.
+NoiseDF [35] raised an innovative idea to distinguish Deepface-specific noise traces from common noises within each single frame, providing additional interpretability.
+With the rapid advancements in vision transformers (ViT), several ViT-based approaches [21, 18, 27, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45] have been developed, showing superior detection performance.
 ICT [21] introduced the Identity Consistency Transformer to capture identity inconsistency within fake faces, while F2Trans designed the High-Frequency Fine-Grained Transformer, achieving more generalized and robust forgery detection.
@@
 The designed modules effectively capture global and local face forgery features more efficiently, without the danger of catastrophic forgetting typical of fully finetuning models [52].
+Figure 4: Overview of the designed face forgery detection framework.
+(a) Overall model structure; (b) Structure of the designed transformer block; (c) Details of the designed adapter layer; (d) Details of the designed LoRA layer.
+Figure 5: Illustration of the Central-Difference Convolution (CDC) pipeline.
 III Proposed Method
@@
 Moreover, $h^{out}$ is obtained by conducting an element-wise summation of the two outputs.
-Figure 3: Overview of the designed face forgery detection framework.
-(a) Overall model structure; (b) Structure of the designed transformer block; (c) Details of the designed adapter layer; (d) Details of the designed LoRA layer.
-Figure 4: Illustration of the Central-Difference Convolution (CDC) pipeline.
+Figure 6: Illustration of forgery style mixture pipeline.
+The designed feature extractor processes real faces and forgery source domain data, and the extracted features are subsequently delivered to the forgery-style mixture module.
 III-B Forgery-aware PEFT
@@
 The pipeline of the CDC operator is illustrated in Fig.
-4.
+5.
 The output feature map $x^{out}$ can be calculated by:
 $x^{out}=\sum_{i\in\Omega}w_{i}(x^{in}_{i}-x_{c}),$ (3)
-Figure 5: Illustration of forgery style mixture pipeline.
-The designed feature extractor processes real faces and forgery source domain data, and the extracted features are subsequently delivered to the forgery-style mixture module.
 where $x_{c}$ indicates the center pixel of the window $\Omega$, and $x_{5}$ represents $x_{c}$ in Fig.
-4.
+5.
 $x_{i}$ indicates the peripheral pixels within $\Omega$.
@@
 $h^{out}=W_{q/k/v}h^{in}+{\rm LoRA}(h^{in}).$ (4)
+Figure 7: Pipeline of the proposed forgery feature mixture process.
 The $W_{q/k/v}$ parameter denotes the pre-trained ImageNet weights fixed during the training process. Moreover, the LoRA module consists of two low-rank matrices:
@@
 In this work, we set the hidden state dimension $d$ of ViT-B to 768, and the rank $r$ is defaulted to 8.
-Figure 6: Pipeline of the proposed forgery feature mixture process.
-The designed forgery-aware PEFT modules bear the following advantages for face forgery detection: (1) they effectively preserve knowledge learned from ImageNet, while adaptively learning forgery-specific information; and (2) they simultaneously capture expressive global features and extract informative local forgery clues, thereby enhancing the final detection performance.
-These advantages jointly contribute to a more generalized approach to face forgery detection.
+The designed LoRA and CDC adapter respectively capture expressive global features and extract informative local forgery clues, thereby enhancing the final detection performance.
+The designed forgery-aware PEFT modules bear the following advantages for face forgery detection: (1) By optimizing a small subset of parameters, our method reduces computational and memory requirements compared to full fine-tuning while also accelerating training, making frequent updates more practical in open-set environments.
+Moreover, storing the backbone weights in the cloud while keeping only a minimal set of parameters on edge devices facilitates deployment on resource-constrained mobile platforms.
+(2) As model sizes continue to grow, our forgery-aware PEFT strategy remains scalable for large models, and the designed modules can be seamlessly integrated with various backbones in a plug-and-play fashion.
+(3) Our approach efficiently incorporates forgery-specific knowledge while preserving the pretrained model’s general knowledge, effectively mitigating overfitting to forgery data and enhancing generalization performance.
 III-C Forgery style mixture
@@
 In this vein, we denote the training dataset as $D^{S}=\{R^{S},F_{1}^{S},F_{2}^{S},...,F_{N}^{S}\}$.
-$R^{S}$ denotes the source real faces, and $F_{i}^{S}$ is the $i_{th}$ source forgery domain, $i\in\{1,2,...,N\}$.
-As such, the testing data can be denoted as $D^{T}=\{R^{T},F_{1}^{T},F_{2}^{T},...,F_{M}^{T}\}$.
-Similarly, $R^{T}$ and $F^{T}$ are target real and forgery domains, while $M$ denotes the number of target forgery domains.
-Our preliminary experiment, illustrated in Fig.
-1, reveals a high false negative rate (FNR) and a low false positive rate (FPR) in cross-domain evaluations, suggesting a negligible domain gap between $R^{S}$ and $R^{T}$ and significant domain gaps between $F^{S}$ and $F^{T}$.
-As such, we need to augment source forgery domains to mitigate this issue.
+$R^{S}$ represents the real source faces, while $F_{i}^{S}$ denotes the $i_{th}$ source forgery domain, where $i\in\{1,2,...,N\}$.
+Following previous methods in face forgery detection, we use FF++ as the training set, which includes multiple training forgery domains (i.e., manipulation types).
+As such, the target testing domain can be denoted as $D^{T}=\{R^{T},F^{T}\}$.
+Similarly, $R^{T}$ and $F^{T}$ are target real and forgery domains.
+Our preliminary experiments, illustrated in Fig.
+1 and Fig.
+2, reveals a high false negative rate (FNR) and a low false positive rate (FPR) in cross-domain evaluations, suggesting a negligible domain gap between $R^{S}$ and $R^{T}$ and significant domain gaps between $F^{S}$ and $F^{T}$.
+As such, we propose to augment source forgery domains to mitigate this issue.
 Fig.
 6 depicts the proposed forgery-style mixture pipeline.
-The training batch $X^{S}=\{R^{S},F_{1}^{S},F_{2}^{S},...,F_{N}^{S}\}$ is processed by our designed feature extractor, which shares weights across all source domains.
-The extracted features from the source forgery domains are combined in the designed forgery-style mixture module, and the mixed features are forwarded to the classifier to perform forgery detection.
+Each training batch, composed of one real domain and multiple fake domains, is processed by our proposed feature extractor, which shares weights across all source domains.
+The extracted features from the source forgery domains are passed to the designed forgery-style mixture module, and the mixed features are subsequently forwarded to the classifier to perform forgery detection.
 Next, we elaborate on the details of the devised forgery style mixture module, as depicted in Fig.
-6.
-This module processes a training batch of real and forged face images sampled from $N$ source domains, where $N=3$, for instance, is shown in Fig.
-6.
+7.
 Green squares denote real faces in this representation, while red circles indicate the forgery faces.
-The input feature $f^{in}$ proceeds to our designed Forgery Style Mixture module.
-We first sort $f^{in}$ to $f^{sort}$ according to the domain labels, where $f_{R}$ represents real features, and $f_{Fi}$ with $i\in\{1,2,3\}$, denotes features from distinct forgery domains.
-The superscript of each sample indicates the feature index of the domain.
-To augment the forgery style, we first shuffle the fake features of $f^{sort}$, ensuring that each forgery domain label of the shuffled $\widetilde{f}^{sort}$ differs from the original $f^{sort}$.
+This module processes each training batch consisting of real and fake face images sampled from the real source domain and $N$ forgery source domains.
+In Fig.
+7, the subscripts of $f$ represent the domain labels, while the superscripts denote the sample index within each domain.
+We illustrate the case of NN=3 forgery domains in Fig.
+7, with 3 samples per domain.
+The features $\mathcal{F}_{in}$ of each input batch is proceeded to our designed Forgery Style Mixture module.
+We first sort $\mathcal{F}_{in}$ to $\mathcal{F}_{sort}$ according to the domain labels, where $f_{R}$ represents real features, and $f_{Fi}$ with $i\in\{1,2,3\}$, denotes features from different forgery domains.
+To augment the forgery style, we first shuffle the fake features of $\mathcal{F}_{sort}$, ensuring that each forgery domain label of the shuffled $\widetilde{\mathcal{F}}_{sort}$ differs from the original $\mathcal{F}_{sort}$.
 In addition, we keep the real feature order unchanged.
-Subsequently, we blend the forgery feature styles of the original $f^{sort}$ and the shuffled $\widetilde{f}^{sort}$.
+Subsequently, we mix the original $\mathcal{F}_{sort}$ and the shuffled $\widetilde{\mathcal{F}}_{sort}$ by fusing their feature statistics.
+The details of this process are provided next.
 We draw inspiration from Adaptive Instance Normalization (AdaIN) [56], a widely employed image style transfer technique. The source style of $p$ can be transferred to the target style of $q$ by:
@@
 where $\mu(p)$ and $\sigma(p)$ are the spatial mean and standard deviation of the feature tensor $p$.
-Inspired by the AdaIN, which utilizes the feature statistics of a target style input to transfer the image style, we first calculate the mixed feature statistics of $f$ and $\tilde{f}$:
-$\gamma_{mix}=\delta\sigma(f)+(1-\delta)\sigma(\tilde{f}),$ (7)
-$\eta_{mix}=\delta\mu(f)+(1-\delta)\mu(\tilde{f}),$ (8)
+Inspired by the AdaIN, which utilizes the feature statistics of a target style input to transfer the image style, we first calculate the mixed feature statistics of $\mathcal{F}_{sort}$ and $\widetilde{\mathcal{F}}_{sort}$:
+[…]
+[…]
 where $\delta$ is a weight sampled from Beta distribution, i.e., $\delta\sim Beta(0.1,0.1)$.
-$\gamma_{mix}$ and $\eta_{mix}$ are the weighted summation of the forgery style statistics of $f$ and $\tilde{f}$.
-Then, the style mixture is performed by the mixed feature statistics $\gamma_{mix}$ and $\eta_{mix}$ to the original feature $f$:
-$f_{mix}={\rm MixStyle}(f)=\gamma_{mix}\frac{f-\mu(f)}{\sigma(f)}+\eta_{mix}.$ (9)
+$\gamma_{mix}$ and $\eta_{mix}$ are the weighted summation of the forgery style statistics of $\mathcal{F}_{sort}$ and $\widetilde{\mathcal{F}}_{sort}$.
+Then, the style mixture is performed by the mixed feature statistics $\gamma_{mix}$ and $\eta_{mix}$ to the original feature $\mathcal{F}_{sort}$:
+[…]
+Finally, we restore the original index order of the mixed features $\mathcal{F}_{mix^{\prime}}$ to obtain $\mathcal{F}_{mix}$, which is forwarded to the classifier for final decision-making.
 Similar to [57], we set a default probability of 0.5 to activate the forgery-style mixture module.
-Finally, we restore the original index order of the mixed features $f^{sort}$ to obtain $f^{out}$, which is forwarded to the classifier for final decision-making.
 In the inference stage, the forgery-style mixture is deactivated.
-During training, the designed forgery-style mixture module augments the source forgery domains, effectively mitigating the overfitting problem and enhance the model’s generalizability.
-This design enhances the model’s generalization for open-set face forgery detection.
+As such, the designed module augments the source forgery domains in the training phase, thereby mitigating the overfitting problem and enhance the model’s generalizability.
+This design effectively improves the model’s ability to generalize in open-set face forgery detection.
 TABLE I: High quality (c23) cross-manipulation detection performance on unseen forgery methods. The trainable parameter number of each model is reported to evaluate the training efficiency.
@@
 CFM [63] TIFS24 25.37M 0.880 0.617 0.814 0.557 0.631 0.516 0.644 0.508 0.742 0.550
+DCL [16] AAAI22 19.35M 0.949 0.877 0.829 0.684 - - - - - -
+Implicit ID [64] CVPR23 22.36M 0.950 0.882 0.844 0.694 - - - - - -
+UCF [65] CVPR23 20.81M - 0.867 - 0.679 - - - - - -
+DiffusionFake [66] NIPS24 19.34M - 0.882 - 0.702 - - - - - -
 SCLoRA [27] MIPR23 4.72M 0.935 0.862 0.875 0.753 0.651 0.554 0.707 0.626 0.792 0.699
@@
 CFM [63] TIFS24 25.37M 0.789 0.686 0.622 0.541 0.688 0.607 0.566 0.517 0.666 0.588
+DCL [16] AAAI22 19.35M 0.838 0.759 0.751 0.679 - - - - - -
+Implicit ID [64] CVPR23 22.40M 0.846 0.769 0.748 0.680 - - - - - -
+UCF [65] CVPR23 20.81M - 0.746 - 0.673 - - - - - -
+DiffusionFake [66] NIPS24 19.34M - 0.751 - 0.713 - - - - - -
 SCLoRA [27] MIPR23 4.72M 0.818 0.735 0.686 0.638 0.710 0.653 0.582 0.542 0.699 0.642
@@
 where $Dist_{R}$ and $Dist_{F}$ denote the average distance between the real feature center $C$ and real feature $f_{R}$ and fake feature $f_{F}$, respectively:
-$Dist_{R}=\frac{1}{N_{R}}\sum_{{i=1}}^{N_{R}}||f_{R}^{i}-C||_{2};\ Dist_{F}=\frac{1}{N_{F}}\sum_{{j=1}}^{N_{F}}||f_{F}^{j}-C||_{2}$ (12)
+[…]
 where the real center $C$ is the average of all $N_{R}$ real features $f_{R}$:
@@
 IV-A1 Implementation details
-In this study, we employ the popular dlib face detector [47] to crop face regions.
+In this study, we employ the popular dlib face detector [67].
+The detected face regions are enlarged by a factor of 1.3 and subsequently resized to 224×224 pixels before being input into the model.
 The proposed framework is implemented on the PyTorch [68] platform, and the model is trained using Adam optimizer [69] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999.
 We set the learning rate at 3e-5, and no learning rate decay is applied throughout the whole training process.
-Our default architecture employs the ViT-B as the backbone, and we train the model on a single 2080Ti GPU with a batch size of 48.
-The training comprises 30,000 iterations, and the ViT-B parameters are initialized with ImageNet-21k [50] weights.
+Our default architecture employs the ViT-B as the backbone, and the ViT-B parameters are initialized with ImageNet-21k [70] weights.
+We set the batch size to 48, as a single 2080Ti GPU can accommodate a maximum of 48 images at a resolution of $\times224$.
+During training, we use AUC as the validation metric.
+The training process consists of 30,000 iterations, as we observe that the validation AUC converges after about 20,000 iterations.
 IV-A2 Datasets
@@
 EN-B4†[12] ICML19 19.34M 65.24 39.41 67.89 37.21 67.96 37.60 63.48 40.14 92.18 15.51 70.03 35.40 71.13 35.40
+Implicit ID [64] CVPR23 22.36M 83.80 24.85 - - - - 81.23 26.80 - - - - - -
+SFDG [77] CVPR23 22.40M 75.83 30.30 69.27 37.70 - - 73.64 33.70 92.10 15.10 - - - -
+UCF* [65] CVPR23 46.80M 77.15 30.20 77.40 29.80 69.30 35.80 68.40 37.80 80.80 27.00 68.30 36.70 73.56 32.88
+LSDA [78] CVPR24 20.81M 82.70 - 71.40 - 76.30 - 71.10 - 87.50 - 84.20 - 78.87 -
+DiffusionFake [66] NIPS24 19.34M 83.17 24.59 75.17 33.25 77.35 30.17 - - - - - - - -
+DSRL* [79] IJCV24 128.31M 74.23 30.48 74.18 31.16 72.01 34.03 73.74 31.35 94.50 13.09 86.77 20.41 79.24 26.75
+FreqNet† [80] AAAI24 1.85M 72.81 33.01 73.37 32.67 72.89 33.37 68.78 36.65 86.94 26.48 75.67 30.26 75.08 32.07
 CFM [63] TIFS24 25.37M 82.78 24.74 78.39 30.79 75.82 31.67 68.28 37.14 95.18 11.87 78.71 29.28 79.86 27.58
@@
 RECCE†[76] CVPR22 25.83M 69.25 34.38 76.99 30.49 66.90 39.39 70.26 34.25 97.15 9.29 92.05 17.74 78.77 27.59
+UCF* [65] CVPR23 46.80M 83.70 28.20 77.20 26.10 71.00 33.50 70.90 35.80 86.40 20.40 69.70 34.20 76.48 29.70
+LSDA [78] CVPR24 20.81M 89.80 - 75.60 - 81.20 - 73.50 - 89.20 - 84.20 - 82.25 -
+DSRL* [79] IJCV24 128.31M 84.56 22.90 77.31 27.32 75.64 30.17 77.99 27.63 97.23 10.44 88.69 19.82 83.57 23.05
+FreqNet†[80] AAAI24 1.85M 75.79 28.59 79.54 29.14 75.07 31.16 69.51 36.11 88.79 25.97 79.56 26.47 78.04 29.57
 CFM [63] TIFS24 25.37M 89.65 17.65 82.27 26.80 80.22 27.48 70.59 35.02 97.59 9.04 83.81 24.49 84.02 23.41
@@
 IV-C Cross-dataset evaluation
-Due to the significant domain gaps between training and testing datasets [59, 60] in face forgery detection, cross-dataset evaluation has been a persistent challenge.
+Due to the significant domain gaps between training and testing datasets [83] in face forgery detection, cross-dataset evaluation has been a persistent challenge.
 We further conduct cross-destatet evaluations to assess the generalizability of forgery detectors.
-We train the models on the FF++ (c23) training set (OR, DF, FF, FS, and NT) and test them on six popular Deepfake datasets, including CelebDF-v2 (CDF) [61], WildDeepfake (WDF) [62], DeepFake Detection Challenge (DFDC) [63], DeepFake Detection Challenge Preview (DFDC-P) [63], DeepForensics-1.0 (DFR) [64], and Face Forensics in the Wild (FFIW) [65].
+To ensure fair comparisons in cross-dataset experiments, we follow previous works [63, 76, 16] by using the FF++ (c23) dataset [22] as the training set.
+This dataset includes original real faces and four types of forgeries: Deepfakes, Face2Face, FaceSwap, and NeuralTextures.
+To balance positive and negative samples during training, real faces are augmented fourfold.
+Additionally, during the forgery style mixture stage, we randomly mix the four forgery types.
+We test the trained models on six popular Deepfake datasets, including CelebDF-v2 (CDF) [84], WildDeepfake (WDF) [25], DeepFake Detection Challenge (DFDC) [23], DeepFake Detection Challenge Preview (DFDC-P) [23], DeepForensics-1.0 (DFR) [24], and Face Forensics in the Wild (FFIW) [26].
 These testing datasets simulate real-world scenarios well, encompassing various environmental variables.
@@
 IV-D Robustness evaluation
-Videos circulating on social media platforms often undergo various distortions during transmission.
-To simulate real-world scenarios, we introduce six common perturbations: Gaussian blur, pink noise, white noise, block, brightness, and darkness.
-Fig.
-8 showcases both the original and corresponding distorted face images.
 We incorporate five severity levels for each perturbation type to mimic unpredictable environments.
 The robustness evaluation results are depicted in Fig.
-7.
+8.
 OSDFD does not employ data augmentations during training, facilitating efficient training, or fine-tuning for real-world applications.
@@
 Furthermore, the average robustness results across all distortions highlight that OSDFD, with the fewest trainable parameters, demonstrates the best robustness against practical image/video perturbations.
+Our proposed style mixture module contributes to enhancing model robustness by increasing the diversity of forgery domains, thereby reducing the distribution gap between clean training data and corrupted testing data.
 Figure 8: Robustness to common perturbations at five severity levels: Gaussian Blur, Pink Noise, White Noise, Block, Brightness, and Darkness.
-Figure 8: Input face image and its perturbated versions.
 TABLE IV: Ablation Experiments on different model components. Average Cross-Manipulation detection performance on FF++ c23 dataset.
@@
 - - - 0.825 0.723
+- - - 0.795 0.712
 - - 0.832 0.740
@@
 Overall, the experimental results presented in Table IV underscore the effectiveness of the designed components.
-The proposed method showcases its adaptability by being flexibly adapted to other ViT backbones, a practical feature for researchers, engineers, and practitioners in the field of computer vision and image processing.
-Herein, we assess the face forgery detection performance of our OSDFD on various ViT backbones, including ViT-Tiny (ViT-T), ViT-Small (ViT-S), and ViT-Large (ViT-L).
-The cross-manipulation (c23) AUC results are presented in Table V.
-It can be readily observed that the average cross-manipulation performance improves with increasing model capacity.
-Nevertheless, larger ViT models tend to require significantly more computing resources and inference time, thereby impeding their real-world deployment.
-In contrast, OSDFD consistently enhances the generalizability of ViT backbones across different model complexities, underscoring its practicality and adaptability.
-Specifically, our method achieves average AUC enhancements of 11.7%, 15.7%, and 16.7%, respectively.
-These results demonstrate the capability of OSDFD to seamlessly integrate into existing ViT models in a plug-and-play manner.
+The proposed method showcases its adaptability by being flexibly adapted to other transformer backbones.
+OSDFD comprises forgery-aware, parameter-efficient fine-tuning modules and a forgery-style mixture module, both designed as plug-and-play components that improve the generalization capability of transformer backbones.
+As shown in Table V, we apply our method OSDFD to various models, including ViT [17] and Swin [85].
+The results demonstrate that OSDFD consistently enhances performance across all backbones, leading to significant improvements in AUC scores in the cross-manipulation setting.
+Notably, the highest performance gains are observed in ViT-L (+0.167 AVG AUC) and Swin-L (+0.112 AVG AUC), suggesting that OSDFD is particularly effective when integrated with larger backbones.
+While the average cross-manipulation performance improves with increasing model capacity, larger models tend to require significantly more computing resources and inference time, thereby impeding their real-world deployment.
+These findings indicate that OSDFD is a flexible and effective module for improving the generalizability of transformer-based face forgery detection models.
 In this work, we adopt a Parameter-Efficient Fine-Tuning (PEFT) strategy, a reliable approach, to train the face forgery detection model.
@@
 Furthermore, the PEFT strategy effectively preserves the abundant ImageNet prior knowledge while learning the forgery-aware knowledge through the designed Adapter and LoRA modules in an adaptive manner, reinforcing the reliability of our research.
-TABLE V: Cross-Manipulation AUC performance on FF++ (c23) dataset with different ViT backbones.
-Backbone OSDFD DF FF FS NT AVG
+TABLE V: Cross-manipulation AUC performance on FF++ (c23) dataset with different backbones.
+Backbone OSDFD DF $F$ FS NT AVG
 ViT-T - 0.853 0.658 0.488 0.584 0.646
@@
 ViT-L 0.955 0.941 0.710 0.785 0.848
+Swin-S - 0.897 0.768 0.400 0.666 0.683
+Swin-S 0.922 0.867 0.538 0.759 0.772
+Swin-B - 0.877 0.715 0.463 0.668 0.681
+Swin-B 0.943 0.888 0.583 0.734 0.787
+Swin-L - 0.923 0.743 0.446 0.646 0.690
+Swin-L 0.956 0.906 0.601 0.745 0.802
 TABLE VI: Cross-Manipulation AUC performance on FF++ (c23) dataset with different training strategies.
-Strategy DF FF FS NT AVG
+Strategy DF $F$ FS NT AVG
 Fully FT 0.896 0.731 0.496 0.603 0.682
@@
 Figure 9: Average cross-manipulation ACC scores across four experimental trials with respect to different mixed feature ratios.
+Figure 10: Cross-dataset evaluation results using different seeds: (a) AUC; (b) EER.
 Figure 11: Grad-CAM [86] visualization results of the ViT-B baseline and OSDFD on six unforeseen Deepfake datasets.
-IV-E2 Impacts of different ratios of mixed features
+IV-E2 Effectiveness of FSM
+To further evaluate the effectiveness of the FSM module, we report additional metrics, including mAP, ACC, and FNR, in Table VII.
+The cross-dataset evaluation results with and without FSM are presented in Table VII, showing that the inclusion of FSM consistently improves face forgery detection performance.
+Notably, FSM significantly enhances the FNR, indicating that it effectively mitigates forgery domain gaps between training and testing sets.
+This, in turn, enhances the model’s ability to detect forgery faces from another point of view.
+IV-E3 Effectiveness of CDC
+To validate the effectiveness of the CDC operator, we conduct an ablation study by integrating various conventional convolutions with different kernel sizes.
+Table VIII presents the model’s cross-dataset experimental results using different local feature extractors, where Conv3×3, Conv5×5, and Conv7×7 represent standard convolutions with kernel sizes of $3\times 3$, $5\times 5$, and $7\times 7$, respectively.
+The Second-order Difference Convolution (SDC) module, as proposed by Fei et al.
+[54], employs a second-order local anomaly learning mechanism to detect anomalies in local regions.
+These baseline results are obtained by replacing the CDC convolution in Fig.
+4 with the corresponding operators.
+The results demonstrate that the CDC operator consistently outperforms other local feature extractors across various unforeseen datasets, which can be attributed to its outstanding capability for local high-frequency feature extraction.
+IV-E4 Impacts of different ratios of mixed features
 Our proposed method utilizes a feature mixture module to augment forgery domains in the feature space.
@@
 In Fig.
-11, we present the t-SNE [67] feature embedding visualization results of both the ViT-B baseline and our method OSDFD.
+12, we present the t-SNE [87] feature embedding visualization results of both the ViT-B baseline and our method OSDFD.
 Fig.
@@
 In contrast, OSDFD effectively separates real and fake images in the feature space.
-This improvement underscores the enhanced generalizability achieved by our model, attributed to the designed forgery-style mixture module and forgery-aware PEFT modules.
+Specifically, for CDF and FFIW datasets, it is obvious that OSDFD can better discriminate real faces from fake ones compared to the ViT baseline.
+The AUC score goes from 72.35 to 83.35 on the CDF dataset, and 68.49 to 84.61 on the FFIW dataset.
+In turn, ViT-B can achieve good detection performance on WDF and DFDC datasets, while our proposed method OSDFD, benefiting from the designed forgery-aware parameter efficient fine-tuning and the forgery style mixture modules, exhibits a better face forgery detection performance.
 Figure 12: t-SNE feature embedding visualization of (a) the ViT-B baseline on the CDF dataset, (b) OSDFD on the CDF dataset, (c) the ViT-B baseline on the WDF dataset, (d) OSDFD on the WDF dataset, (e) the ViT-B baseline on the DFDC dataset, (f) OSDFD on the DFDC dataset, (g) the ViT-B baseline on the FFIW dataset, and (h) OSDFD on the FFIW dataset.
 The red and blue dots represent real and fake data samples, respectively.
+TABLE VII: Impacts of Forgery Style Mixture (FSM).
+Setting CDF DFR WDF
+[…]
+w/o FSM 88.91 72.87 30.69 93.42 84.28 15.40 84.74 70.94 23.61
+w/ FSM 90.33 75.05 22.82 95.25 88.22 11.01 86.70 72.77 16.13
 IV-F2 Grad-CAM map visualization
 We visualize the Grad-CAM [86] maps to identify attention regions of the developed face forgery detector for input queries, as shown in Fig.
-10.
+11.
 These maps illustrate the results of the ViT-B baseline model and OSDFD across six unforeseen Deepfake datasets: CDF, WDF, DFDC, DFDC-P, DFR, and FFIW.
@@
 These observations further underscore the outstanding generalizability of our method in real-world applications.
+TABLE VIII: Comparison with other local-frequency extractor.
+Setting CDF DFR WDF AVG
+[…]
+Conv3×3 82.26 26.66 94.57 11.59 78.00 29.20 84.94 22.48
+Conv5×5 80.39 27.32 91.35 15.75 76.98 30.18 82.91 24.42
+Conv7×7 80.44 27.71 92.92 14.25 78.10 29.09 83.82 23.68
+SDC [54] 79.14 27.89 92.87 11.21 77.72 28.32 83.24 22.47
+CDC 83.35 24.47 95.54 9.90 78.53 28.46 85.81 20.94
+IV-F3 Cross-dataset results using different seeds
+To mitigate the influence of different seeds, we train our models for seven trials using seeds 0-6.
+The cross-dataset AUC and EER performances are shown in Fig.
+10 (a) and (b), respectively.
+The results indicate that our approach consistently achieves promising performance regardless of the different seeds, demonstrating its robustness across varied initializations.
+This robustness further underscores the superiority of our method from an additional perspective.
 V Conclusions and Future Work
@@
 Adapting our approach to leverage abundant sequential information could significantly improve video-level face forgery detection.
-The designed components can be seamlessly integrated into video transformers, achieving superior performance in uncontrolled environments.
-This path can be explored in future work.
+The designed components can be seamlessly integrated into video transformers.
+Extending our designed method to video transformers for video-level Deepfake detection opens an important research path forward.
+We anticipate that the proposed modules can well adapt to video ViTs and enhance the generalizability.
+Furthermore, our method exhibits strong robustness against common image perturb ations, as shown in Fig.
+8.
+However, its performance degrades under the highest perturbation severity (level 5).
+Additionally, although our approach contributes to safeguarding digital media integrity, Deepfake detection models may exhibit bias across different demographic groups due to imbalances in the training dataset.
+In future work, our aim is to improve robustness and ensure fairness in Deepfake detection.

### v3

@@
 Open-set face forgery detection poses significant security threats and presents substantial challenges for existing detection models.
-These detectors primarily have two limitations: they cannot generalize across unknown forgery domains and inefficiently adapt to new data.
+These detectors primarily have two limitations: they cannot generalize across unknown forgery domains or inefficiently adapt to new data.
 To address these issues, we introduce an approach that is both general and parameter-efficient for face forgery detection.
 Our method builds on the assumption that different forgery source domains exhibit distinct style statistics.
-Specifically, we design a forgery-style mixture formulation that augments the diversity of forgery source domains, enhancing the model’s generalizability across unseen domains.
+Specifically, we design a forgery-style-mixture formulation that augments the diversity of forgery source domains, enhancing the model’s generalizability across unseen domains.
 In addition, previous methods typically require fully fine-tuning pretrained networks, consuming substantial time and computational resources.
 Drawing on recent advancements in vision transformers (ViT) for face forgery detection, we develop a parameter-efficient ViT-based detection model that includes lightweight forgery feature extraction modules and enables the model to extract global and local forgery clues simultaneously.
-We only optimize the inserted lightweight modules during training, maintaining the original ViT structure with its pre-trained ImageNet weights.
+We only optimize the inserted lightweight modules during training, maintaining the original ViT structure with its pre-trained weights.
 This training strategy effectively preserves the informative pre-trained knowledge while flexibly adapting the model to the task of Deepfake detection.
@@
 Later approaches, including Face X-ray [14], SBI [15], DCL [16], and RECCE [16] sought to capture common forgery features to enhance model generalizability.
-With the advent of ViT [17], numerous ViT-based detectors [18, 19, 20, 21] have been developed.
+With the advent of ViT [17], numerous ViT-based detectors [18, 19, 20, 21, 22] have been developed.
 These approaches exploit global receptive fields and exhibit superior detection performance.
@@
 Motivated by this observation, we design the forgery style mixture module to enhance the diversity of source forgery domains, thereby improving the model’s generalizability.
-Furthermore, most prior art utilizes generic CNNs and ViTs as backbones and fully fine-tune the networks from ImageNet weights.
+Furthermore, most prior art utilizes generic CNNs and ViTs as backbones and fully fine-tune the networks from weights.
 However, fully fine-tuning a network is costly, hindering its real-world applications.
 To mitigate this issue, we integrate lightweight Adapter and LoRA layers into the ViT backbone.
-During training, we freeze plain ViT backbones with ImageNet weights and solely update the lightweight Adapter and LoRA layers.
+During training, we freeze plain ViT backbones with weights and solely update the lightweight Adapter and LoRA layers.
 These designed modules capture local and global forgery artifacts using slim components.
 Additionally, these modules preserve substantial ImageNet knowledge and enable flexible model adaptation for face forgery detection.
-While parameter-efficient fine-tuning (PEFT) has been widely adopted in various domains, our work makes a notable contribution by integrating PEFT with a forgery-aware perspective, specifically tailored for open-set face forgery detection—a task that presents unique challenges in generalizability and adaptation.
+While parameter-efficient fine-tuning (PEFT) has been widely adopted in various domains, our work makes a contribution by integrating PEFT with a forgery-aware perspective, specifically tailored for open-set face forgery detection—a task that presents unique challenges in generalizability and adaptation.
 Fig.
@@
 In summary, the main contributions of this work are as follows:
-  • We design a forgery-style mixture module to enhance the diversity of source domains. The model’s generalizability is effectively improved by randomly mixing forgery styles.
+  • We design a forgery-style-mixture module to enhance the diversity of source domains. The model’s generalizability is effectively improved by randomly mixing forgery styles.
   • Our approach explicitly captures forgery-related artifacts by integrating a CDC adapter and a LoRA module into a ViT-based detection framework.
@@
   • Our designed components jointly contribute to an effective and scalable solution for open-set Deepfake detection.
-Extensive experiments demonstrate that our model achieves state-of-the-art generalizability and robustness with only 1.34M activated parameters.
+Extensive experiments demonstrate that our model achieves state-of-the-art generalizability and robustness with the fewest activated parameters.
 We arranged the remaining of this paper into four more sections, Sec.
@@
 With the rise of deep learning and artificial intelligence, numerous learning-based methods emerged to improve detection accuracy.
-Some generic CNNs, such as the Xception network [11, 29, 30, 7, 31], EfficientNet-b4 [12, 32, 33, 34], and Capsule network [13], have been utilized for face forgery detection, showing promising results.
+Some generic CNNs, such as the Xception network [11, 30, 31, 7, 32], EfficientNet-b4 [12, 33, 34, 35, 36], and Capsule network [13], have been utilized for face forgery detection, showing promising results.
 However, these methods often need to be more balanced in training data, leading to poor generalization capabilities.
@@
 NoiseDF [37] raised an innovative idea to distinguish Deepface-specific noise traces from common noises within each single frame, providing additional interpretability.
-With the rapid advancements in vision transformers (ViT), several ViT-based approaches [21, 18, 27, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45] have been developed, showing superior detection performance.
-ICT [21] introduced the Identity Consistency Transformer to capture identity inconsistency within fake faces, while F2Trans designed the High-Frequency Fine-Grained Transformer, achieving more generalized and robust forgery detection.
+With the rapid advancements in vision transformers (ViT), several ViT-based approaches [21, 18, 28, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 39] have been developed, showing superior detection performance.
+ICT [21] introduced the Identity Consistency Transformer to capture identity inconsistency within fake faces, while F2Trans [48] designed the High-Frequency Fine-Grained Transformer, achieving more generalized and robust forgery detection.
+Forensics Adapter [49] is a lightweight network that equips CLIP with forgery-specific knowledge by learning blending boundary traces and enhancing visual tokens through a dedicated interaction strategy.
+TFCU [50] improves face forgery video detection by sequentially capturing momentary anomalies, gradual inconsistencies, and cumulative distortions through its consecutive correlation, future-guided, and historical review modules.
+OWG-DS [51] advances real-world Deepfake detection by transferring knowledge from limited labeled data to large-scale unlabeled data.
+GM-DF [52] enhances Deepfake detection generalization across multiple datasets, enabling effective handling of diverse real-world scenarios and unseen attacks.
+DAID [53] mitigates demographic biases in Deepfake detection through demographic-aware data rebalancing and demographic-agnostic feature aggregation.
+Recent LVLM-based approaches [54, 55, 56, 57, 58, 59, 60] take a further step forward by achieving explainable face forgery detection.
 However, the scalability of ViT-based methods is largely constrained by the quadratic complexity of attention computation, limiting their deployment, particularly on mobile devices with limited computing resources.
@@
 Despite these advancements, the development of PEFT methods for face forgery detection has been largely understudied, yet it is crucial in practical scenarios.
-In this paper, we integrate lightweight Adapter and LoRA modules into the plain ViT backbone, optimizing only these modules while keeping the ViT backbone fixed with ImageNet weights during training.
+In this paper, we integrate lightweight Adapter and LoRA modules into the plain ViT backbone, optimizing only these modules while keeping the ViT backbone fixed with pretrained weights during training.
 The designed modules effectively capture global and local face forgery features more efficiently, without the danger of catastrophic forgetting typical of fully finetuning models [67].
-Figure 4: Overview of the designed face forgery detection framework.
-(a) Overall model structure; (b) Structure of the designed transformer block; (c) Details of the designed adapter layer; (d) Details of the designed LoRA layer.
-Figure 5: Illustration of the Central-Difference Convolution (CDC) pipeline.
 III Proposed Method
@@
 Moreover, $h^{out}$ is obtained by conducting an element-wise summation of the two outputs.
-Figure 6: Illustration of forgery style mixture pipeline.
-The designed feature extractor processes real faces and forgery source domain data, and the extracted features are subsequently delivered to the forgery-style mixture module.
 III-B Forgery-aware PEFT
@@
 ${\rm Adapter}(h^{in})={\rm Conv_{1\times 1}^{up}}({\rm CDC}({\rm Conv_{1\times 1}^{down}}(h^{in}))).$ (2)
+Figure 4: Overview of the designed face forgery detection framework.
+(a) Overall model structure; (b) Structure of the designed transformer block; (c) Details of the designed adapter layer; (d) Details of the designed LoRA layer.
+Figure 5: Illustration of the Central-Difference Convolution (CDC) pipeline.
 We first expand the input hidden state $h^{in}$ into a 2-D dimension.
@@
 Finally, the second $1\times 1$ Conv layer restores the channels of the feature, and the produced feature is flattened to the same shape as $h^{in}$.
+CDC operates within a local sliding window, calculating differences between peripheral and central pixels to extract high-frequency features [68].
+The CDC adapter functions as a local anomaly extractor for the ViT backbone by capturing subtle forgery cues, such as boundary inconsistencies and local irregularities.
+These local cues are then integrated with the backbone’s global features, while the PEFT strategy ensures this integration is achieved effectively with minimal trainable parameters.
 Our designed adapter layer effectively introduces inductive bias into standard ViTs, guiding the model in learning local artifacts for face forgery detection.
@@
 4 (d), the output hidden state $h^{out}$ is calculated by:
+Figure 6: Illustration of forgery style mixture pipeline.
+The designed feature extractor processes real faces and forgery source domain data, and the extracted features are subsequently delivered to the forgery-style-mixture module.
 $h^{out}=W_{q/k/v}h^{in}+{\rm LoRA}(h^{in}).$ (4)
-Figure 7: Pipeline of the proposed forgery feature mixture process.
-The $W_{q/k/v}$ parameter denotes the pre-trained ImageNet weights fixed during the training process. Moreover, the LoRA module consists of two low-rank matrices:
+The $W_{q/k/v}$ parameter denotes the pre-trained weights fixed during the training process. Moreover, the LoRA module consists of two low-rank matrices:
 ${\rm LoRA}(h^{in})=h^{in}W_{q/k/v}^{down}W_{q/k/v}^{up},$ (5)
@@
 Moreover, storing the backbone weights in the cloud while keeping only a minimal set of parameters on edge devices facilitates deployment on resource-constrained mobile platforms.
-(2) As model sizes continue to grow, our forgery-aware PEFT strategy remains scalable for large models, and the designed modules can be seamlessly integrated with various backbones in a plug-and-play fashion.
+(2) As model sizes continue to grow, our forgery-aware PEFT strategy remains scalable for large models, and the designed modules can be seamlessly integrated with various transformer-based backbones in a plug-and-play fashion.
 (3) Our approach efficiently incorporates forgery-specific knowledge while preserving the pretrained model’s general knowledge, effectively mitigating overfitting to forgery data and enhancing generalization performance.
+Figure 7: Pipeline of the proposed forgery feature mixture process.
 III-C Forgery style mixture
@@
 Fig.
-6 depicts the proposed forgery-style mixture pipeline.
+6 depicts the proposed forgery-style-mixture pipeline.
 Each training batch, composed of one real domain and multiple fake domains, is processed by our proposed feature extractor, which shares weights across all source domains.
-The extracted features from the source forgery domains are passed to the designed forgery-style mixture module, and the mixed features are subsequently forwarded to the classifier to perform forgery detection.
+The extracted features from the source forgery domains are passed to the designed forgery-style-mixture module, and the mixed features are subsequently forwarded to the classifier to perform forgery detection.
 Next, we elaborate on the details of the devised forgery style mixture module, as depicted in Fig.
@@
 Finally, we restore the original index order of the mixed features $\mathcal{F}_{mix^{\prime}}$ to obtain $\mathcal{F}_{mix}$, which is forwarded to the classifier for final decision-making.
-Similar to [57], we set a default probability of 0.5 to activate the forgery-style mixture module.
-In the inference stage, the forgery-style mixture is deactivated.
+Similar to [73], we set a default probability of 0.5 to activate the forgery-style-mixture module.
+In the inference stage, the forgery-style-mixture is deactivated.
 As such, the designed module augments the source forgery domains in the training phase, thereby mitigating the overfitting problem and enhance the model’s generalizability.
 This design effectively improves the model’s ability to generalize in open-set face forgery detection.
-TABLE I: High quality (c23) cross-manipulation detection performance on unseen forgery methods. The trainable parameter number of each model is reported to evaluate the training efficiency.
+In contrast to conventional computer vision tasks, forgery detection suffers from imbalanced domain gaps: the discrepancies between different fake domains are substantially larger than those among real domains, due to the continual evolution of manipulation techniques.
+Forgery Style Mixture (FSM) is a label-guided mechanism that explicitly mixes feature statistics across different forgery styles.
+Unlike prior augmentation strategies that enhance general feature diversity, FSM specifically targets forgery-style diversity, thereby expanding the fake feature space without requiring additional data.
+TABLE I: High quality (c23) cross-manipulation detection performance. (#Params: trainable parameter number)
 […]
@@
 LTW [78] AAAI21 20.37M 0.927 0.856 0.802 0.656 0.640 0.549 0.773 0.653 0.786 0.679
-ViT-B [17] ICLR21 85.80M 0.771 0.701 0.656 0.582 0.510 0.498 0.554 0.517 0.623 0.575
 CFM [79] TIFS24 25.37M 0.880 0.617 0.814 0.557 0.631 0.516 0.644 0.508 0.742 0.550
@@
 SCLoRA [28] MIPR23 4.72M 0.935 0.862 0.875 0.753 0.651 0.554 0.707 0.626 0.792 0.699
-OSDFD Ours 1.34M 0.962 0.885 0.932 0.846 0.698 0.595 0.781 0.682 0.843 0.752
-TABLE II: Low quality (c40) cross-manipulation detection performance on unseen forgery methods. The trainable parameter number of each model is reported to evaluate the training efficiency.
+ViT-B [17] ICLR21 85.80M 0.771 0.701 0.656 0.582 0.510 0.498 0.554 0.517 0.623 0.575
+CLIP [83] ICML21 303.18M 0.890 0.766 0.847 0.737 0.541 0.504 0.652 0.538 0.733 0.636
+OSDFD (ViT-B) Ours 1.34M 0.962 0.885 0.932 0.846 0.698 0.595 0.781 0.682 0.843 0.752
+OSDFD (CLIP) Ours 2.89M 0.990 0.934 0.945 0.858 0.881 0.790 0.793 0.703 0.902 0.821
+TABLE II: Low quality (c40) cross-manipulation detection performance. (#Params: trainable parameter number)
 […]
@@
 LTW [78] AAAI21 20.37M 0.756 0.691 0.724 0.657 0.681 0.625 0.608 0.585 0.692 0.640
-ViT-B [17] ICLR21 85.80M 0.739 0.643 0.650 0.595 0.592 0.560 0.552 0.538 0.633 0.584
 CFM [79] TIFS24 25.37M 0.789 0.686 0.622 0.541 0.688 0.607 0.566 0.517 0.666 0.588
@@
 SCLoRA [28] MIPR23 4.72M 0.818 0.735 0.686 0.638 0.710 0.653 0.582 0.542 0.699 0.642
-OSDFD Ours 1.34M 0.853 0.762 0.717 0.663 0.713 0.661 0.630 0.595 0.728 0.670
+ViT-B [17] ICLR21 85.80M 0.739 0.643 0.650 0.595 0.592 0.560 0.552 0.538 0.633 0.584
+CLIP [83] ICML21 303.18M 0.800 0.707 0.676 0.615 0.651 0.606 0.592 0.544 0.680 0.618
+OSDFD (ViT-B) Ours 1.34M 0.853 0.762 0.717 0.663 0.713 0.661 0.630 0.595 0.728 0.670
+OSDFD (CLIP) Ours 2.89M 0.924 0.819 0.761 0.684 0.817 0.716 0.628 0.601 0.783 0.705
 III-D Objective function
@@
 $L=L_{BCE}+\lambda L_{SCL},$ (10)
-where $\lambda$ balances the two loss components.
+$\lambda$ balances the two loss components.
 $L_{BCE}$ represents the binary cross-entropy loss.
@@
 $L_{SCL}=Dist_{R}+max(Dist_{R}-Dist_{F}+margin,0)$ (11)
-where $Dist_{R}$ and $Dist_{F}$ denote the average distance between the real feature center $C$ and real feature $f_{R}$ and fake feature $f_{F}$, respectively:
+where margin (default: 0.01) is a hyperparameter that specifies how much farther the fake features should be from the real center compared to the real features.
+$Dist_{R}$ and $Dist_{F}$ denote the average distance between the real feature center $C$ and real feature $f_{R}$ and fake feature $f_{F}$, respectively:
 […]
@@
 This loss enhances the discriminative and separable characteristics of the real and fake face features, thereby improving the overall performance of face forgery detection.
+FSM is coupled with Single Center Loss ($L_{SCL}$) to jointly form a forgery-oriented regularization strategy.
 IV Experiments and Results
@@
 We set the learning rate at 3e-5, and no learning rate decay is applied throughout the whole training process.
-Our default architecture employs the ViT-B as the backbone, and the ViT-B parameters are initialized with ImageNet-21k [70] weights.
-We set the batch size to 48, as a single 2080Ti GPU can accommodate a maximum of 48 images at a resolution of $\times224$.
+In the proposed OSDFD framework, the ViT-Base model pre-trained on ImageNet-21K [87] serves as the backbone.
+It consists of 12 transformer blocks, each comprising a 12-head self-attention layer and a multi-layer perceptron (MLP) layer.
+We use the ViT-L/14 as the CLIP backbone.
+We set the batch size to 48, and the loss weight $\lambda$ in Eq.
+(10) is set to 1.
 During training, we use AUC as the validation metric.
@@
 Unless specified otherwise, we report the image-level detection results throughout this paper.
-IV-B Cross-manipulation evaluation
-In real-world applications, face forgery detectors inevitably encounter data generated by new forgery techniques, leading to significant performance drops.
-To evaluate the generalization capability, we train the models on the training sets comprising three source manipulation types of FF++.
-Subsequently, we evaluate the trained models on the test sets of the unseen target manipulation type.
-Table I presents the detection AUC and ACC of different models on the FF++ high-quality (c23) dataset.
-We bold the best detection results and underline the second-best results.
-It can be observed that OSDFD achieves SOTA AUC and ACC performances across all experiments.
-Moreover, generalization to poor-quality data is crucial for real-world face forgery detectors.
-Therefore, we conduct the cross-manipulation experiments on low-quality (c40) data, as shown in Table II, where OSDFD outperforms previous approaches by a significant margin.
-This underscores the proposed method’s outstanding generalizability across diverse data qualities.
-Additionally, our approach demonstrates substantial improvements over the ViT-B baseline, demonstrating the effectiveness of the designed framework.
-Furthermore, compared to our conference version SCLoRA [27], our method consistently achieves superior performance on both c23 and c40 data, which can be attributed to the effectiveness of the designed forgery style mixture module and local CDC feature extractor.
-Finally, OSDFD comprises the fewest 1.34M trainable parameters, facilitating its deployment in real-world applications, particularly on mobile devices.
 TABLE III: Cross-dataset evaluation on six unseen datasets.
@@
 ‘†’ indicates our re-implementation using the public official code.
-Methods highlighted in blue denote video-level results.
-The trainable parameter number of each model is reported to evaluate the training efficiency.
+Top and bottom tables show frame- and video-level results, respectively.
+The trainable parameter number of each model is also reported.
+We separate the CLIP-based methods from others.
 Method Venue #Params CDF WDF DFDC-P DFDC DFR FFIW Avgerage
@@
 CFM [79] TIFS24 25.37M 82.78 24.74 78.39 30.79 75.82 31.67 68.28 37.14 95.18 11.87 78.71 29.28 79.86 27.58
+LAA-Net* [36] CVPR24 27.12M 81.99 25.34 73.83 32.59 75.71 31.61 66.97 37.71 80.56 26.77 72.38 34.05 75.24 31.34
 ViT-B [17] ICLR21 85.80M 72.35 34.50 75.29 33.40 75.58 32.11 71.48 35.54 80.47 26.97 68.49 37.12 73.94 33.27
 SCLoRA [28] MIPR23 4.72M 73.56 32.87 77.91 29.56 77.08 30.37 72.32 34.52 94.78 11.76 81.42 26.59 79.51 26.59
-OSDFD Ours 1.34M 83.35 24.47 78.53 28.46 76.28 31.17 75.52 31.99 95.54 9.90 84.61 24.09 82.31 25.01
+OSDFD (ViT-B) Ours 1.34M 83.35 24.47 78.53 28.46 76.28 31.17 75.52 31.99 95.54 9.90 84.61 24.09 82.31 25.01
+ForAda* [49] CVPR25 7.54M 87.61 21.71 79.90 27.60 83.87 24.74 79.42 28.08 93.51 13.96 77.66 29.62 83.67 24.29
+UDD* [97] AAAI25 0.66M 82.68 25.20 75.92 31.31 79.56 27.92 73.67 32.42 90.93 16.84 78.79 28.10 80.26 26.97
+CLIP [83] ICML21 303.18M 81.23 26.46 76.41 31.55 77.97 29.76 72.54 33.97 95.16 11.69 83.63 25.19 81.15 24.77
+OSDFD (CLIP) Ours 2.89M 88.32 20.45 85.39 21.97 79.42 30.00 77.88 30.15 99.17 3.46 93.71 14.88 87.32 20.15
 Lisiam [98] TIFS22 - 78.21 - - - - - - - - - - - - -
@@
 CFM [79] TIFS24 25.37M 89.65 17.65 82.27 26.80 80.22 27.48 70.59 35.02 97.59 9.04 83.81 24.49 84.02 23.41
+LAA-Net* [36] CVPR24 27.12M 91.99 15.73 78.84 27.78 78.78 28.72 70.28 35.26 87.86 19.90 77.97 29.75 80.95 26.19
 ViT-B [17] ICLR21 85.80M 78.12 30.59 78.95 29.59 79.43 28.62 73.35 33.05 86.42 19.90 70.34 35.45 77.77 29.53
 SCLoRA [28] MIPR23 4.72M 81.11 26.32 82.78 24.85 80.20 26.68 74.96 35.01 98.53 6.63 85.59 22.70 83.86 23.70
-OSDFD Ours 1.34M 90.81 16.94 81.42 24.90 79.49 28.31 78.74 29.37 98.89 4.48 90.01 17.50 86.56 20.25
+OSDFD (ViT-B) Ours 1.34M 90.81 16.94 81.42 24.90 79.49 28.31 78.74 29.37 98.89 4.48 90.01 17.50 86.56 20.25
+ForAda* [49] CVPR25 7.54M 93.87 14.41 85.64 22.47 91.40 16.53 83.77 23.51 97.48 8.46 85.12 23.22 89.55 18.10
+UDD* [97] AAAI25 0.66M 90.31 17.29 78.90 32.20 84.97 23.76 77.14 29.85 96.17 11.44 85.54 23.51 85.51 23.01
+CLIP [83] ICML21 303.18M 88.87 21.47 80.33 28.54 80.17 27.76 76.07 30.94 98.67 6.97 90.91 17.22 85.83 22.15
+OSDFD (CLIP) Ours 2.89M 93.23 16.29 87.16 19.51 83.13 25.76 80.65 27.53 99.62 1.49 97.20 9.76 90.17 16.72
+IV-B Cross-manipulation evaluation
+In real-world applications, face forgery detectors inevitably encounter data generated by new forgery techniques, leading to significant performance drops.
+To evaluate the generalization capability, we train the models on the training sets comprising three source manipulation types of FF++.
+Subsequently, we evaluate the trained models on the test sets of the unseen target manipulation type.
+Table I presents the detection AUC and ACC of different models on the FF++ high-quality (c23) dataset.
+We bold the best detection results and underline the second-best results.
+We apply OSDFD to two backbone networks, ViT-B and CLIP (ViT-L/14).
+The results show that OSDFD built on ViT-B surpasses previous methods, while OSDFD (CLIP) achieves SOTA AUC and ACC across all evaluation settings.
+Moreover, generalization to poor-quality data is crucial for real-world face forgery detectors.
+Therefore, we conduct the cross-manipulation experiments on low-quality (c40) data, as shown in Table II, where OSDFD outperforms previous approaches by a clear margin under three settings.
+This underscores the proposed method’s outstanding generalizability across diverse data qualities.
+OSDFD exhibits relatively lower performance on the (DF, FS, NT $\rightarrow$ FF) setting under low-quality cross-manipulation evaluation.
+This may be because the PEFT framework and style mixture module are well suited for capturing appearance-related anomalies but are less effective at generalizing to manipulations dominated by expression and motion artifacts, thereby limiting transferability to FF.
+Additionally, our approach demonstrates substantial improvements over the ViT-B baseline, demonstrating the effectiveness of the designed framework.
+Furthermore, compared to our conference version SCLoRA [28], our method achieves superior performance on both c23 and c40 data, which can be attributed to the effectiveness of the designed forgery style mixture module and local CDC feature extractor.
+Moreover, OSDFD (CLIP) exhibits good generalizability on low-quality data while requiring only 2.89M trainable parameters, making it well suited for deployment in open-set scenarios.
+TABLE IV: Cross-dataset evaluations on DF40 (CDF) and FakeAVCeleb datasets.
+Method FS (CDF) FR (CDF) EFS (CDF) FakeAVCeleb Average
+AUC EER AUC EER AUC EER AUC EER AUC EER
+SBI[15] 70.01 35.20 48.90 50.98 65.61 39.21 85.87 23.07 67.59 37.11
+CFM[79] 78.65 28.73 69.11 35.96 71.49 34.33 82.84 26.27 75.52 31.32
+DCL[16] 72.67 35.04 62.88 41.01 63.18 41.20 87.64 21.00 71.59 34.56
+LAA-Net[36] 70.46 34.99 55.24 46.91 67.07 38.07 88.20 19.74 70.24 34.92
+DiffusionFake[82] 70.34 35.75 55.37 46.85 63.70 41.57 85.91 22.54 68.83 36.67
+ViT-B [17] 67.78 37.19 55.65 46.57 60.50 42.64 82.98 26.14 66.72 38.13
+OSDFD (ViT-B) 82.50 24.59 64.78 39.22 77.68 26.65 84.04 23.03 77.25 28.37
+CLIP [83] 78.93 29.21 62.33 42.49 71.23 34.10 82.86 25.27 73.83 32.76
+ForAda[49] 83.33 25.13 63.76 40.73 82.30 25.34 83.67 25.41 78.26 29.15
+UDD[97] 61.44 41.64 59.01 43.63 52.64 47.54 90.72 16.02 65.95 37.20
+OSDFD (CLIP) 90.77 16.97 71.06 35.00 90.09 18.70 92.95 14.54 86.21 21.30
+TABLE V: Efficiency evaluations w/ and w/o OSDFD.
+[…]
+[…]
+[…]
+[…]
+(98.44% $\downarrow$) (9.34% $\uparrow$) (22.49% $\downarrow$) (+1.31) (+0.16) (+8.37)
+[…]
+[…]
+(99.05% $\downarrow$) (20.48% $\uparrow$) (7.91% $\downarrow$) (+3.68) (+0.71) (+6.17)
 IV-C Cross-dataset evaluation
-Due to the significant domain gaps between training and testing datasets [83] in face forgery detection, cross-dataset evaluation has been a persistent challenge.
+Due to the significant domain gaps between training and testing datasets [100, 101] in face forgery detection, cross-dataset evaluation has been a persistent challenge.
 We further conduct cross-destatet evaluations to assess the generalizability of forgery detectors.
-To ensure fair comparisons in cross-dataset experiments, we follow previous works [63, 76, 16] by using the FF++ (c23) dataset [22] as the training set.
-This dataset includes original real faces and four types of forgeries: Deepfakes, Face2Face, FaceSwap, and NeuralTextures.
+In cross-dataset evaluation, our method OSDFD adopts the FF++ (c23) dataset [23] as the training set for fair comparisons.
+Our training set includes original real faces and four types of forgeries: Deepfakes, Face2Face, FaceSwap, and NeuralTextures, and the four manipulation types are randomly mixed during the forgery style mixture.
 To balance positive and negative samples during training, real faces are augmented fourfold.
-Additionally, during the forgery style mixture stage, we randomly mix the four forgery types.
 We test the trained models on six popular Deepfake datasets, including CelebDF-v2 (CDF) [102], WildDeepfake (WDF) [26], DeepFake Detection Challenge (DFDC) [24], DeepFake Detection Challenge Preview (DFDC-P) [24], DeepForensics-1.0 (DFR) [25], and Face Forensics in the Wild (FFIW) [27].
@@
 AUC and EER performance are reported in Table III, where ‘*’ indicates the trained model provided by the authors, and ‘†’ indicates our re-implementation using the public official code.
-The upper table presents the frame-level detection results, while the blue-highlighted bottom table shows video-level results.
+The upper table presents the frame-level detection results, while the bottom table shows video-level results.
 Video-level detection scores are calculated by averaging all frame scores within each video.
+Since CLIP serves as a more powerful backbone for face forgery detection, we separate the CLIP-based methods (ForAda [49], UDD [97], CLIP [83], and OSDFD (CLIP)) from the other approaches to ensure fair comparisons.
 The best and second-best results are highlighted in bold and underlined, respectively.
 The trainable parameter number of each detector is provided in the table for reference.
-Our method, OSDFD, achieves the highest average detection performance in both frame-level and video-level evaluation, with only 1.34M activated parameters.
-Compared to the ViT-B backbone, OSDFD consistently demonstrates significant performance improvements across all six datasets, demonstrating its exceptional generalization capability.
-Furthermore, OSDFD outperforms our conference version SCLoRA by a significant margin, underscoring the effectiveness of the designed CDC Adapter layer and forgery-style mixture module.
+OSDFD (ViT-B) achieves promising average detection performance in both frame-level and video-level evaluations.
+In contrast, OSDFD built on the CLIP backbone attains better AUC and EER, as CLIP’s multimodal pretraining enables it to learn representations that align visual and semantic concepts, thereby enhancing generalization to distribution shifts.
+OSDFD’s relatively lower performance on DFDC and DFDC-P can be attributed to their large domain gaps with FF++ and the suppression of forgery artifacts by compression, which pose significant challenges for PEFT-based methods.
+Compared to the ViT-B and CLIP backbones, OSDFD demonstrates significant performance improvements across all six datasets, demonstrating its exceptional generalization capability.
+Furthermore, OSDFD outperforms our conference version SCLoRA by a significant margin, underscoring the effectiveness of the designed CDC Adapter layer and forgery-style-mixture module.
+We expand our evaluation to two other recently released and challenging datasets: DF40 [103] and FakeAVCeleb [104].
+DF40 encompasses 40 diverse face forgery methods.
+Its domain-specific fake data generated from FF++ and CDF real faces.
+FakeAVCeleb is a multimodal benchmark comprising 20,000 videos, built from real YouTube celebrity footage across four ethnic backgrounds.
+We train our models on FF++ and report frame-level cross-dataset evaluations in Table IV.
+To prevent facial identity leakage, we test on three challenging DF40 subsets (FS (CDF), FR (CDF), and EFS (CDF)) derived from CDF real faces, and additionally evaluate on FakeAVCeleb (visual modality only).
+The upper part of Table IV shows results with conventional backbones, while the lower part reports CLIP-based models.
+Despite the sophistication of the testing datasets, OSDFD (ViT-B) achieves the best average AUC in the upper table, while OSDFD (CLIP) consistently outperforms all compared methods across all subsets, demonstrating superior generalizability on recent challenging datasets.
+The superior performance of CLIP-based models compared to ViT-B arises from CLIP’s large-scale multimodal pretraining, which equips it with stronger generalization ability to unseen forgery patterns.
 IV-D Robustness evaluation
@@
 As the severity of distortion increases, all face forgery detection methods experience performance degradation.
-Compared to previous arts, our proposed approach achieves superior or comparable AUC performance across all six types of perturbation.
-Furthermore, the average robustness results across all distortions highlight that OSDFD, with the fewest trainable parameters, demonstrates the best robustness against practical image/video perturbations.
-Our proposed style mixture module contributes to enhancing model robustness by increasing the diversity of forgery domains, thereby reducing the distribution gap between clean training data and corrupted testing data.
+The average robustness results across the six distortion types indicate that both OSDFD (ViT-B) and OSDFD (CLIP) exhibit superior resilience to common image perturbations.
+OSDFD (ViT) demonstrates a lower robustness than EfficientNet-B4 under brightness perturbations.
+A potential reason lies in our use of the PEFT strategy.
+Pretrained knowledge captures real visual statistics (e.g., natural lighting), whereas forgeries frequently violate these subtle priors (e.g., illumination inconsistencies), leaving detectable forgery cues.
+However, brightness perturbations during robustness evaluation can obscure these artifacts, thereby diminishing our method’s performance.
+The designed FSM module enhances robustness by generating broader feature distributions to mitigate overfitting, introducing variability that improves resistance to perturbations, and guiding the model to focus on intrinsic forgery cues rather than superficial styles.
 Figure 8: Robustness to common perturbations at five severity levels: Gaussian Blur, Pink Noise, White Noise, Block, Brightness, and Darkness.
-TABLE IV: Ablation Experiments on different model components. Average Cross-Manipulation detection performance on FF++ c23 dataset.
+TABLE VI: Ablation Experiments on different model components. Average Cross-Manipulation detection performance on FF++ c23 dataset.
 ViT LoRA Adapter MixStyle SCL AUC ACC
@@
 0.843 0.752
-IV-E Ablation study
+IV-E Efficiency evaluations
+OSDFD demonstrates high training efficiency and strong generalization capability with only negligible additional FLOPs.
+By introducing a small number of trainable parameters, OSDFD achieves lower storage requirements, faster training, and reduced GPU memory consumption compared with the vanilla backbones.
+We conducted comprehensive efficiency evaluations of OSDFD using two backbones, ViT-B and CLIP, as presented in Table V.
+All experiments were implemented on an NVIDIA A40 GPU and an Intel(R) Xeon(R) Gold 6530 CPU.
+The training speed and inference time were measured over five trials, while the training speed and GPU memory were recorded with a batch size of 32.
+For reference, we also report the average AUC (AVG.
+AUC), representing the mean performance across six unseen datasets.
+Compared with the vanilla ViT-B and CLIP backbones, OSDFD demonstrates strong parameter efficiency by introducing only a small number of trainable parameters.
+This design results in lower storage requirements, as the backbone weights can be stored in the cloud while only minimal parameters are retained locally.
+Moreover, OSDFD enables faster training because fewer parameters are updated during optimization.
+As shown in Table V, OSDFD improves training speed by 9.34% and 20.48% over the ViT-B and CLIP baselines, respectively.
+This efficiency makes OSDFD well-suited for frequent model updates, such as continual learning with newly emerging forgery data in open-set Deepfake detection.
+Furthermore, OSDFD effectively reduces GPU memory consumption during training, making it practical for deployment on resource-constrained devices.
+The designed PEFT and forgery-style-mixture components introduce only a marginal increase in model size, leading to negligible additional inference time and computational overhead.
+Specifically, OSDFD incurs only 1.31 ms and 3.68 ms inference time, and introduces merely 0.16 G and 0.71 G additional FLOPs for the ViT-B and CLIP backbones, respectively, which are acceptable for real-world applications.
+Beyond training efficiency, OSDFD provides additional benefits.
+The forgery-aware PEFT and forgery-style-mixture modules effectively integrate forgery-specific knowledge while preserving the pretrained model’s general representations.
+This design mitigates overfitting to forgery data and enhances generalization performance.
+IV-F Ablation study
 In this subsection, we carry out extensive ablation experiments to showcase the effectiveness of the designed components, the flexibility of the proposed learning scheme, the rationale behind the adopted training strategy, and the impacts of the mixed features.
-IV-E1 Effectiveness of the designed modules
+IV-F1 Effectiveness of the designed modules
 We conduct ablation experiments to study the impact of each designed module.
-Table IV reports the average cross-manipulation (c23) detection AUC and ACC scores across four experimental trials.
+Table VI reports the average cross-manipulation (c23) detection AUC and ACC scores across four experimental trials.
 The designed LoRA and Adapter modules significantly enhance the model’s generalizability.
 The combination of LoRA and Adapter enables the model to simultaneously explore global and local anomalies within input forgery face images, enhancing the final detection performance.
-The forgery-style mixture module further boosts the AUC and ACC scores by augmenting the forgery source domains during training.
+The forgery-style-mixture module further boosts the AUC and ACC scores by augmenting the forgery source domains during training.
 The Single-Center Loss (SCL) encourages the real features to cluster more tightly while pushing features from fake images away from the real center.
 This design leads to a clearer decision boundary between real and fake faces, further enhancing the generalizability of the detection.
-Overall, the experimental results presented in Table IV underscore the effectiveness of the designed components.
-The proposed method showcases its adaptability by being flexibly adapted to other transformer backbones.
-OSDFD comprises forgery-aware, parameter-efficient fine-tuning modules and a forgery-style mixture module, both designed as plug-and-play components that improve the generalization capability of transformer backbones.
-As shown in Table V, we apply our method OSDFD to various models, including ViT [17] and Swin [85].
-The results demonstrate that OSDFD consistently enhances performance across all backbones, leading to significant improvements in AUC scores in the cross-manipulation setting.
+Overall, the experimental results presented in Table VI underscore the effectiveness of the designed components.
+The proposed method showcases its adaptability by being flexibly adapted to other transformer-based backbones.
+OSDFD comprises forgery-aware, parameter-efficient fine-tuning modules and a forgery-style-mixture module, both designed as plug-and-play components that improve the generalization capability of transformer backbones.
+As shown in Table VII, we apply our method OSDFD to various models, including ViT [17] and Swin [105].
+The results demonstrate that OSDFD enhances performance across these backbones, leading to significant improvements in AUC scores in the cross-manipulation setting.
 Notably, the highest performance gains are observed in ViT-L (+0.167 AVG AUC) and Swin-L (+0.112 AVG AUC), suggesting that OSDFD is particularly effective when integrated with larger backbones.
@@
 These findings indicate that OSDFD is a flexible and effective module for improving the generalizability of transformer-based face forgery detection models.
-In this work, we adopt a Parameter-Efficient Fine-Tuning (PEFT) strategy, a reliable approach, to train the face forgery detection model.
-Specifically, we only optimize the Adapter and LoRA layers while keeping the other ViT parameters fixed with ImageNet weights.
-We compare the PEFT strategy with the full fine-tuning (FT) strategy to validate its effectiveness.
-Table VI presents the cross-manipulation (c23) AUC results, where DF, FF, FS, and NT denote the testing manipulation types.
-The models are trained on three manipulations and tested on the remaining one.
-The two experimental settings employ two training strategies while maintaining the same network architectures.
-The results in Table VI demonstrate that the PEFT approach outperforms the fully FT strategy across all experimental trials.
-Notably, the average AUC score of PEFT exhibits a substantial enhancement, increasing from 68.2% to 84.3%.
-Hence, it can be concluded that the improvements in face forgery detection performance benefit from the PEFT strategy adopted.
-Furthermore, the PEFT strategy effectively preserves the abundant ImageNet prior knowledge while learning the forgery-aware knowledge through the designed Adapter and LoRA modules in an adaptive manner, reinforcing the reliability of our research.
-TABLE V: Cross-manipulation AUC performance on FF++ (c23) dataset with different backbones.
+TABLE VII: Cross-manipulation AUC performance on FF++ (c23) dataset with different backbones.
 Backbone OSDFD DF $F$ FS NT AVG
@@
 Swin-L 0.956 0.906 0.601 0.745 0.802
-TABLE VI: Cross-Manipulation AUC performance on FF++ (c23) dataset with different training strategies.
+In this work, we adopt a Parameter-Efficient Fine-Tuning (PEFT) strategy, a reliable approach, to train the face forgery detection model.
+Specifically, we only optimize the Adapter and LoRA layers while keeping the other ViT parameters fixed with ImageNet weights.
+We compare the PEFT strategy with the full fine-tuning (FT) strategy to validate its effectiveness.
+Table VIII presents the cross-manipulation (c23) AUC results, where DF, FF, FS, and NT denote the testing manipulation types.
+The models are trained on three manipulations and tested on the remaining one.
+The two experimental settings employ two training strategies while maintaining the same network architectures.
+The results in Table VIII demonstrate that the PEFT approach outperforms the fully FT strategy across all experimental trials.
+Notably, the average AUC score of PEFT exhibits a substantial enhancement, increasing from 68.2% to 84.3%.
+Hence, it can be concluded that the improvements in face forgery detection performance benefit from the PEFT strategy adopted.
+Furthermore, the PEFT strategy effectively preserves the abundant ImageNet prior knowledge while learning the forgery-aware knowledge through the designed Adapter and LoRA modules in an adaptive manner, reinforcing the reliability of our research.
+Although ImageNet is not face-specific, its learned representations of natural visual statistics (e.g., smooth textures, natural lighting) remain highly relevant, as forgeries often violate these subtle priors (e.g., abnormal texture smoothness, illumination mismatches).
+A backbone pretrained on ImageNet therefore provides strong priors regarding what “real” images should look like, facilitating more effective forgery detection.
+Retaining these pretrained weights preserves natural image priors that help distinguish real from fake distributions, while PEFT introduces forgery-aware adaptability without overwriting the pretrained knowledge, thereby improving generalization and robustness to unseen domains.
+Moreover, since forgery artifacts are diverse and continuously evolving techniques, a backbone grounded in natural image distributions offers a stable reference frame, enabling consistent detection across varied manipulation types.
+TABLE VIII: Cross-Manipulation AUC performance on FF++ (c23) dataset with different training strategies.
 Strategy DF $F$ FS NT AVG
@@
 PEFT (Ours) 0.962 0.932 0.698 0.781 0.843
-Figure 9: Average cross-manipulation ACC scores across four experimental trials with respect to different mixed feature ratios.
-Figure 10: Cross-dataset evaluation results using different seeds: (a) AUC; (b) EER.
-Figure 11: Grad-CAM [86] visualization results of the ViT-B baseline and OSDFD on six unforeseen Deepfake datasets.
-IV-E2 Effectiveness of FSM
-To further evaluate the effectiveness of the FSM module, we report additional metrics, including mAP, ACC, and FNR, in Table VII.
-The cross-dataset evaluation results with and without FSM are presented in Table VII, showing that the inclusion of FSM consistently improves face forgery detection performance.
+TABLE IX: Impacts of Forgery Style Mixture (FSM).
+Setting CDF DFR WDF
+[…]
+w/o FSM 88.91 72.87 30.69 93.42 84.28 15.40 84.74 70.94 23.61
+w/ FSM 90.33 75.05 22.82 95.25 88.22 11.01 86.70 72.77 16.13
+IV-F2 Effectiveness of FSM
+To further evaluate the effectiveness of the FSM module, we report additional metrics, including mAP, ACC, and FNR, in Table IX.
+The cross-dataset evaluation results with and without FSM are presented in Table IX, showing that the inclusion of FSM effectively improves face forgery detection performance.
 Notably, FSM significantly enhances the FNR, indicating that it effectively mitigates forgery domain gaps between training and testing sets.
 This, in turn, enhances the model’s ability to detect forgery faces from another point of view.
-IV-E3 Effectiveness of CDC
+IV-F3 Impacts of training forgery type number
+To investigate the effect of the number of training forgery types, we conducted ablation experiments under the cross-dataset setting.
+As shown in Table X, without the Forgery Style Mixture (FSM) module, incorporating more forgery types improves cross-dataset performance, which aligns with the expectation that larger and more diverse training sets enhance generalizability.
+In contrast, when the FSM module is applied, the model achieves further performance gains on unseen datasets, further confirming its effectiveness in diversifying training forgery style statistics.
+Moreover, the effectiveness of FSM increases with the number of forgery types.
+These findings suggest a promising direction for incorporating additional forgery domains to further enrich forgery style diversity and enhance generalizability.
+Figure 9: Average cross-manipulation ACC scores across four experimental trials with respect to different mixed feature ratios.
+Figure 10: Grad-CAM [106] visualization results of the ViT-B baseline and OSDFD on six unforeseen Deepfake datasets.
+IV-F4 Effectiveness of CDC
 To validate the effectiveness of the CDC operator, we conduct an ablation study by integrating various conventional convolutions with different kernel sizes.
-Table VIII presents the model’s cross-dataset experimental results using different local feature extractors, where Conv3×3, Conv5×5, and Conv7×7 represent standard convolutions with kernel sizes of $3\times 3$, $5\times 5$, and $7\times 7$, respectively.
+Table XI presents the model’s cross-dataset experimental results using different local feature extractors, where Conv3×3, Conv5×5, and Conv7×7 represent standard convolutions with kernel sizes of $3\times 3$, $5\times 5$, and $7\times 7$, respectively.
 The Second-order Difference Convolution (SDC) module, as proposed by Fei et al.
@@
 4 with the corresponding operators.
-The results demonstrate that the CDC operator consistently outperforms other local feature extractors across various unforeseen datasets, which can be attributed to its outstanding capability for local high-frequency feature extraction.
-IV-E4 Impacts of different ratios of mixed features
+The results demonstrate that the CDC operator outperforms other local feature extractors across various unforeseen datasets, which can be attributed to its outstanding capability for local high-frequency feature extraction.
+IV-F5 Impacts of different $\lambda$ values
+To investigate the effect of the loss weight of $L_{SCL}$, we conducted ablation experiments under the cross-dataset setting, as reported in Table XII.
+Compared with the baseline ($\lambda$=0), incorporating the Single-Center Loss $L_{SCL}$ effectively enhances the model’s generalizability.
+We varied the $\lambda$ value from 0.6 to 1.4 and found that $\lambda$=1.0 achieves better performance across unseen datasets.
+Figure 11: T-SNE visualizations on FF++ and CDF datasets.
+(a) ViT-B baseline; (b) ViT-B with Forgery Style Mixture (FSM) module.
+(DF: Deepfakes, $F$: Face2Face, FS: FaceSwap, NT: NeuralTextures, OR: Original.)
+IV-F6 Impacts of different ratios of mixed features
 Our proposed method utilizes a feature mixture module to augment forgery domains in the feature space.
@@
 Similarly, we present the average score across four cross-manipulation evaluation trials.
-A ratio of ‘0.0’ indicates the pure use of mixed forgery features during training, while ‘1.0’ means no forgery-style mixture is applied.
+A ratio of ‘0.0’ indicates the pure use of mixed forgery features during training, while ‘1.0’ means no forgery-style-mixture is applied.
 Notably, relying solely on original or mixed forgery features results in limited face forgery detection accuracy.
@@
 This observation underscores the effectiveness of the designed forgery feature mixture module from another perspective.
-IV-F Visualization results
+TABLE X: Ablation experiment on training manipulation type number (cross-dataset evaluation).
+FSM DF $F$ FS NT CDF DFR WDF
+AUC EER AUC EER AUC EER
+- - - - 75.85 32.08 91.22 15.86 72.13 34.43
+- - - 77.51 29.80 92.13 15.61 73.68 33.81
+- - 78.03 29.71 92.44 14.78 74.26 31.91
+- - 78.12 30.01 93.01 14.57 75.18 31.82
+- 79.88 28.26 93.77 12.88 76.75 30.89
+- 80.76 26.58 94.18 11.69 76.68 30.76
+83.35 24.47 95.54 9.90 78.53 28.46
+IV-G Visualization results
 In this subsection, we provide visualization results to illustrate further the effectiveness of the proposed face forgery detection method.
-IV-F1 T-SNE feature map visualization
+IV-G1 T-SNE feature map visualization
+We visualize the embedding distributions of ViT-B baseline (a) w/o and (b) w/ the Forgery Style Mixture (FSM) module in Fig.
+11 to demonstrate the effectiveness of FSM, which is a key component of our method.
+The models are trained on the FF++ training sets and tested on both the FF++ test sets and the unseen CDF dataset.
+Feature centers of each domain are marked with $\star$”.
+In Fig.
+11(a), the FF++ (OR) and CDF (Real) domains are closely aligned, whereas the unseen CDF (Fake) domain is significantly distant from the FF++ fake domains.
+This explains why the model effectively detects real faces but struggles with fake ones.
+Motivated by this observation, we design a FSM module to diversify the fake training data and expand the fake feature space without introducing additional data.
+As shown in Fig.
+11(b), the FSM module effectively enlarges the forgery feature space, leading to greater overlap between the FF++ fake domains and the unseen CDF (Fake) domain.
+Consequently, the decision boundary becomes more robust, enabling improved detection of unseen face forgeries.
 In Fig.
@@
 In turn, ViT-B can achieve good detection performance on WDF and DFDC datasets, while our proposed method OSDFD, benefiting from the designed forgery-aware parameter efficient fine-tuning and the forgery style mixture modules, exhibits a better face forgery detection performance.
-Figure 12: t-SNE feature embedding visualization of (a) the ViT-B baseline on the CDF dataset, (b) OSDFD on the CDF dataset, (c) the ViT-B baseline on the WDF dataset, (d) OSDFD on the WDF dataset, (e) the ViT-B baseline on the DFDC dataset, (f) OSDFD on the DFDC dataset, (g) the ViT-B baseline on the FFIW dataset, and (h) OSDFD on the FFIW dataset.
-The red and blue dots represent real and fake data samples, respectively.
-TABLE VII: Impacts of Forgery Style Mixture (FSM).
-Setting CDF DFR WDF
-[…]
-w/o FSM 88.91 72.87 30.69 93.42 84.28 15.40 84.74 70.94 23.61
-w/ FSM 90.33 75.05 22.82 95.25 88.22 11.01 86.70 72.77 16.13
-IV-F2 Grad-CAM map visualization
+IV-G2 Grad-CAM map visualization
 We visualize the Grad-CAM [106] maps to identify attention regions of the developed face forgery detector for input queries, as shown in Fig.
-11.
+10.
 These maps illustrate the results of the ViT-B baseline model and OSDFD across six unforeseen Deepfake datasets: CDF, WDF, DFDC, DFDC-P, DFR, and FFIW.
@@
 Conversely, our method effectively identifies regions with significant forgery artifacts, contributing to superior face forgery detection performance.
-Specifically, our approach consistently demonstrates outstanding gradient-based localization across diverse environmental variables, such as extreme head poses in WDF, abnormal color hues in DFDC, facial occlusion (glasses) in DFDC-P, and extreme facial area ratios in FFIW.
+Specifically, our approach demonstrates an outstanding gradient-based localization across diverse environmental variables, such as extreme head poses in WDF, abnormal color hues in DFDC, facial occlusion (glasses) in DFDC-P, and extreme facial area ratios in FFIW.
 These observations further underscore the outstanding generalizability of our method in real-world applications.
-TABLE VIII: Comparison with other local-frequency extractor.
+Figure 12: T-SNE feature embedding visualization of (a) the ViT-B baseline on the CDF dataset, (b) OSDFD on the CDF dataset, (c) the ViT-B baseline on the WDF dataset, (d) OSDFD on the WDF dataset, (e) the ViT-B baseline on the DFDC dataset, (f) OSDFD on the DFDC dataset, (g) the ViT-B baseline on the FFIW dataset, and (h) OSDFD on the FFIW dataset.
+The red and blue dots represent real and fake data samples, respectively.
+TABLE XI: Comparison with other local-frequency extractor.
 Setting CDF DFR WDF AVG
@@
 CDC 83.35 24.47 95.54 9.90 78.53 28.46 85.81 20.94
-IV-F3 Cross-dataset results using different seeds
+TABLE XII: Ablation experiment on $\lambda$ (cross-dataset evaluation).
+$\lambda$ CDF DFR WDF
+AUC EER AUC EER AUC EER
+0 79.85 28.86 93.62 14.04 76.82 31.71
+0.6 81.33 26.98 94.88 12.69 78.21 30.81
+0.8 82.20 25.43 95.28 11.66 79.03 28.31
+1.0 83.35 24.47 95.54 9.90 78.53 28.46
+1.2 81.93 25.52 95.02 12.51 78.44 29.02
+1.4 80.86 24.41 94.66 11.96 77.97 31.05
+TABLE XIII: Impacts of different seeds.
+CDF DFDC-P DFR WDF FFIW
+[…]
+[…]
+IV-G3 Cross-dataset results using different seeds
 To mitigate the influence of different seeds, we train our models for seven trials using seeds 0-6.
-The cross-dataset AUC and EER performances are shown in Fig.
-10 (a) and (b), respectively.
-The results indicate that our approach consistently achieves promising performance regardless of the different seeds, demonstrating its robustness across varied initializations.
+The cross-dataset AUC and EER performances are shown in Table.
+XIII.
+The results indicate that our approach achieves promising performance regardless of the different seeds, demonstrating its robustness across varied initializations.
 This robustness further underscores the superiority of our method from an additional perspective.
@@
 Our method, OSDFD, incorporates lightweight LoRA and Adapter layers into ViT backbones, updating only these modules during training to learn global and local forgery-aware features efficiently.
-On the other hand, we propose a novel forgery-style mixture module that augments the forgery source domains.
+On the other hand, we propose a novel forgery-style-mixture module that augments the forgery source domains.
 In this vein, our method achieves SOTA generalizability and robustness in face forgery detection.
 An ablation study further underscores that the designed network modules and objective functions contribute to the final detection performance.
-Our proposed method, OSDFD, can be flexibly adapted to other backbones in a plug-and-play manner.
+Our proposed method, OSDFD, can be flexibly adapted to other transformer-based backbones in a plug-and-play manner.
 Further ablation experiments demonstrate that the adopted parameter-efficient training strategy is effective in face forgery detection.
@@
 In future work, our aim is to improve robustness and ensure fairness in Deepfake detection.
+VI Acknowledgement
+This research is supported in part by the National Nature Science Foundation of China (NSFC) under Grant 62502187; in part by the Natural Science Foundation of Jiangxi Province of China under Grant 20252BAC240015; in part by Sichuan Science and Technology Fund 2025ZNSFSC0511; in part by the Research Grant Council (RGC) of Hong Kong SAR, under a GRF Grant 12203124 and an ECS Grant 22201125; in part by the São Paulo Research Foundation (Fapesp) Horus project, Grant #2023/12865-8 and the Brazilian National Council for Scientific and Technological Development (CNPq), Aletheia project, Grant #442229/2024-0.
+This work was carried out at the Rapid-Rich Object Search (ROSE) Lab, School of Electrical & Electronic Engineering, Nanyang Technological University (NTU), Singapore.
+This research is supported in part by A*STAR under its OTS Research Programme (Award S24T2TS006).
+Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the A*STAR.